### v1

# DeepProtein: Deep Learning Library and Benchmark for Protein Sequence Learning

Abstract
In recent years, deep learning has revolutionized the field of protein science, enabling advancements in predicting protein properties, structural folding and interactions. This paper presents DeepProtein, a comprehensive and user-friendly deep learning library specifically designed for protein-related tasks. DeepProtein integrates a couple of state-of-the-art neural network architectures, which include convolutional neural network (CNN), recurrent neural network (RNN), transformer, graph neural network (GNN), and graph transformer (GT). It provides user-friendly interfaces, facilitating domain researchers in applying deep learning techniques to protein data. Also, we curate a benchmark that evaluates these neural architectures on a variety of protein tasks, including protein function prediction, protein localization prediction, and protein-protein interaction prediction, showcasing its superior performance and scalability. Additionally, we provide detailed documentation and tutorials to promote accessibility and encourage reproducible research. This library is extended from a well-known drug discovery library, DeepPurpose and publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
1 Introduction
Understanding the representation of proteomics is vital in developing traditional biological and medical progress (Wu et al., 2022b; Fu et al., 2024), multi-omics genomics (Wu et al., 2022a; Chen et al., 2021), and curing human diseases (Chen et al., 2024c, b). Being the working house of the cell, it provides lots of functions that support human’s daily life, such as catalyzing biochemical reactions that happen in the body as a role of enzymes and providing helpful immune responses against detrimental substance that acts as immunoglobulin (Wu et al., 2024). Under the necessity of analyzing those useful proteins, several related protein databases are available to researchers (Berman et al., 2000; Bairoch and Apweiler, 2000; Consortium, 2015; Pontén et al., 2008). Apart from the 2D database, some recent 3D Protein Database used AlphaFold 2.0 (Jumper et al., 2021) is important to better assist in learning those representations in 3d-dimensional space. The success of AlphaFold 2.0 has sparked a significant increase in interest in using machine learning techniques for protein learning tasks, of which the goal is to improve our understanding of proteins’ biochemical mechanisms.
Deep learning has shown its power in protein learning tasks, including protein-protein interaction (Gainza et al., 2020), protein folding (Jumper et al., 2021; Lu, 2022; Panou and Reczko, 2020; Chen et al., 2016), protein-ligand interaction (Li et al., 2021b; Xia et al., 2023), and protein function and property prediction (Gligorijević et al., 2021; Sevgen et al., 2023). Convolutional neural network (CNN) (Shanehsazzadeh et al., 2020) and TAPE Transformer (Rao et al., 2019) on proteins are models with regard to protein sequence-to-sequence (Seq2Seq) learning, including the pretrained transformer such as ProtBert (Brandes et al., 2022). With the help of 3D information intrinsics, graph neural networks (GNN) are applied, especially some pretrained GNN models (Jing et al., 2020; Zhang et al., 2022), and graph transformers (Yuan et al., 2022; Gu et al., 2023). While transformer claims to be the state-of-the-art in the previous benchmark (Xu et al., 2022), the analysis of potential powerful deep learning models such as GNN models and graph transformer models are neglected, which prompts us to integrate those methods in our benchmark. Also, pretaining strategies are well-known for its ability in utilizing unlabeled data (Lu et al., 2024; Yue et al., 2024).
Challenges. Previous benchmarks related to molecular learning have offered valuable insights regarding their respective libraries and implementation interfaces. DeepPurpose [^1] https://github.com/kexinhuang12345/DeepPurpose (Huang et al., 2020) has provided an interface that implements the task with a majority of drug discovery tasks, which only has protein-protein interaction and protein function prediction implemented. Datasets on proteins are lacking as well. TorchProtein [^2] https://github.com/DeepGraphLearning/PEER_Benchmark  (Xu et al., 2022), also named as PEER, implemented most of the tasks in the protein field. In terms of models, the focus has largely been on sequential learning methods: Convolutional Neural Networks (CNNs), Transformers, and ESM architectures. This suggests that there are still many structural methods available for consideration. Furthermore, PEER’s interface is not user-friendly without prior domain knowledge in graphs and biochemistry. This situation presents an opportunity to improve the existing interface with regard to simplicity and comprehensibility.
Solutions. To address these challenges, in this paper, we propose DeepProtein, which aims to benchmark mainstream and cutting-edge deep learning models on a wide range of AI-solvable protein sequence learning tasks. We investigate the performance of various deep learning models on a wide range of protein sequence learning tasks. We analyze each method’s advantages and disadvantages when performing each task (working as the explainer for each task). We have provided user-friendly and well-wrapped interfaces to facilitate domain experts’ research.
Contribution. The major contributions of our paper are summarized as
  • Comprehensive Benchmarking: We curate a benchmark to evaluate the performance of 8 neural network architectures (including CNNs, RNNs, transformers, and various graph neural networks) on these 7 essential protein learning tasks (including protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, antibody paratope prediction, CRISPR repair outcome prediction, and antibody developability prediciton), demonstrating superior performance and scalability.
  • User-friendly Library: We develop DeepProtein, a specialized deep learning library that integrates these neural network architectures (CNNs, RNNs, transformers, and graph neural networks) on these 7 protein-related tasks such as protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, and antibody developability prediciton. For each task, our library enables one command line to run every method.
  • Enhanced Accessibility: We provide extensive documentation and tutorials to facilitate user engagement and promote reproducible research, building upon the foundation of the well-known drug discovery library, DeepPurpose (Huang et al., 2020).
    Figure 1: DeepProtein framework.
2 Related Works
Benchmarks and libraries are crucial in AI-based therapeutic science, e.g., multi-omics data (Lu, 2018), protein learning (Xu et al., 2022), small-molecule drug discovery (Gao et al., 2022; Zheng et al., 2024; Xu et al., 2024), and drug development (clinical trial) (Chen et al., 2024a; Wang et al., 2024). They provide standardized metrics for evaluating the performance of various algorithms and models. These benchmarks enable researchers to compare different approaches systematically, ensuring reproducibility and reliability of results.
In this section, we briefly discuss the benchmark studies in this area. Proteins are vital in drug discovery because they often serve as the primary targets for therapeutic agents, influencing disease mechanisms and biological pathways. Additionally, proteins play key roles in various cellular processes, making them essential for identifying potential drug candidates and biomarkers in the drug development pipeline. A couple of protein learning benchmarks are developed, including PEER (Xu et al., 2022), DeepPurpose (Huang et al., 2020), FLIP (Dallago et al., 2021), TAPE (Rao et al., 2019). Table 1 compares DeepProtein with existing AI-based protein learning benchmarks. We extend the scope of existing protein learning benchmarks by incorporating more protein learning datasets, more cutting-edge deep learning models, and enhancing user-friendliness.
[…]
Datasets DeepPurpose FLIP TAPE PEER TDC (data only) DeepProtein
References (Huang et al., 2020) (Dallago et al., 2021) (Rao et al., 2019) (Xu et al., 2022) (Huang et al., 2021) ours
Fluorescence $\times$ $\times$
$\beta$-lactamase $\times$ $\times$ $\times$ $\times$
Solubility $\times$ $\times$ $\times$ $\times$
Stability $\times$ $\times$
Subcellular (Binary) $\times$ $\times$ $\times$ $\times$
PPI Affinity $\times$ $\times$ $\times$ $\times$
Yeast PPI $\times$ $\times$ $\times$ $\times$
Human PPI $\times$ $\times$ $\times$ $\times$
[…]
user-friendly $\times$ $\times$ $\times$
3 DeepProtein Library and Benchmark
3.1 AI-solvable Protein Problems
In this section, we elaborate on a couple of AI-solvable protein problems and the related datasets.
  • Protein Function Prediction. Protein function prediction involves determining the biological roles and activities of proteins based on their sequences or structures. This process is crucial for understanding cellular mechanisms and interactions, as a protein’s function is often linked to its sequence composition and the context of its cellular environment. Machine learning algorithms are employed to analyze known protein databases, identifying patterns and features that correlate with specific functions. Accurate predictions can facilitate drug discovery, help elucidate disease mechanisms, and support advancements in synthetic biology by providing insights into how proteins can be engineered for desired activities (Zhang et al., 2021). We consider the following datasets.
    • Fluorescence (Sarkisyan et al., 2016). Protein fluorescence refers to the phenomenon where certain proteins can emit light of a specific wavelength when excited by light of a shorter wavelength. It is a widely used technique to study protein structure, dynamics, interactions, and function. The dataset consists of 54,025 protein sequences with real-valued groundtruth. The label is the logarithm of fluorescence intensity.
    • Stability (Rocklin et al., 2017). Protein stability is the capacity of a protein to preserve its three-dimensional structure and functional characteristics across different environmental conditions. This stability is essential for the proper functioning and longevity of proteins within biological systems. A protein’s stability is influenced by its ability to withstand denaturation, aggregation, and degradation. The dataset comprises 68,934 protein sequences with real-valued groundtruth.
    • $\beta$-lactamase (Gray et al., 2018). This task aims to predict the increased activity of $\beta$-lactamase, the most common enzyme that provides gram-negative bacteria with resistance to beta-lactam antibiotics through single mutations. The dataset consists of 5,198 protein sequences with real-valued groundtruth. The groundtruth refers to the experimentally determined fitness score, which measures the scaled mutation effect for each mutant.
    • Solubility (Khurana et al., 2018). Protein solubility is the capacity of a protein to dissolve or remain dispersed in a solution. This property is crucial for determining how the protein behaves and functions in various biological and industrial contexts. Several factors influence a protein’s solubility, including its amino acid composition, ionic strength, pH, temperature, and the presence of other molecules in the solution. The dataset consists of 71,419 protein sequences with binary labels.
  • Protein Localization Prediction. Accurate localization predictions can enhance drug development by informing target identification and improving therapeutic efficacy, particularly in treating diseases linked to protein mislocalization. Additionally, insights gained from localization predictions facilitate the mapping of biological pathways, aiding in the identification of new therapeutic targets and potential disease mechanisms.
    • Subcellular (Almagro Armenteros et al., 2017). The task predicts the location of a natural protein within the cell. The dataset consists of 13,961 data samples with categorical labels (10 classes, $\{0,1,2,\cdots,9\}$).
    • Binary (Almagro Armenteros et al., 2017). It is a simpler version of the previous task (10-category classification), where the model is trained to roughly forecast each protein as either “membrane-bound” or “soluble” (i.e., binary classification). The dataset comprises 8,634 data samples with binary labels.
  • Protein-Protein Interaction (PPI). Proteins are the essential functional units in human biology, but they seldom operate in isolation; rather, they typically interact with one another to perform various functions. Understanding protein-protein interactions (PPIs) is crucial for identifying potential therapeutic targets for disease treatment. Traditionally, determining PPI activity requires costly and time-consuming wet-lab experiments. PPI prediction seeks to forecast the activity of these interactions based on the amino acid sequences of paired proteins.
    • PPI Affinity (Moal and Fernández-Recio, 2012). It consists of 2,682 protein-protein pairs with real-valued groundtruth.
    • Yeast (Guo et al., 2008). The dataset comprises 2,172 protein-protein pairs with binary labels.
    • Human PPI (Pan et al., 2010). The dataset comprises 7,348 protein-protein pairs with binary labels.
  • Epitope Prediction. An epitope, also known as an antigenic determinant, is the region of a pathogen that can be recognized by antibodies and cause an adaptive immune response. The epitope prediction task is to distinguish the active and non-active sites from the antigen protein sequences. Identifying the potential epitope is of primary importance in many clinical and biotechnologies, such as vaccine design and antibody development, and for our general understanding of the immune system (Du et al., 2023). In epitope prediction, the machine learning model makes a binary prediction for each amino acid residue. This is also known as residue-level classification.
    • Immune Epitope Database (IEDB) (Vita et al., 2019). It consists of 3,159 antigens with binary labels on each amino acid. The label indicates whether the amino acid belongs to the epitope, i.e., active position in binding. It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/epitope/).
    • PDB-Jespersen (Jespersen et al., 2017). It consists of 447 antigens with binary labels on each amino acid. It is curated by (Jespersen et al., 2017) and is extracted from PDB (Protein Data Bank). It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/epitope/).
  • Paratope Prediction. Antibodies, or immunoglobulins, are large, Y-shaped proteins that can recognize and neutralize specific molecules on pathogens, known as antigens. They are crucial components of the immune system and serve as valuable tools in research and diagnostics. The paratope, also referred to as the antigen-binding site, is the region that specifically binds to the epitope. While we have a general understanding of the hypervariable regions responsible for this binding, accurately identifying the specific amino acids involved remains a challenge. This task focuses on predicting which amino acids occupy the active positions of the antibody that interact with the antigen. In paratope prediction, the machine learning model makes a binary prediction for each amino acid residue. This is also known as residue-level classification.
    • SAbDab-Liberis (Liberis et al., 2018) is curated from SAbDab (Dunbar et al., 2014). It consists of 1,023 antibody chain sequences; each antibody contains both heavy and light chain sequences. It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/paratope/#sabdab-liberis-et-al).
  • Antibody Developability Prediction. Immunogenicity, instability, self-association, high viscosity, polyspecificity, and poor expression can hinder an antibody from being developed as a therapeutic agent, making early identification of these issues crucial. The goal of antibody developability prediction is to predict an antibody’s developability from its amino acid sequences. A fast and reliable developability predictor can streamline antibody development by minimizing the need for wet lab experiments, alerting chemists to potential efficacy and safety concerns, and guiding necessary modifications. While previous methods have used 3D structures to create accurate developability indices, acquiring 3D information is costly. Therefore, a machine learning approach that calculates developability based solely on sequence data is highly advantageous.
    • TAP (Raybould et al., 2019). It contains 242 antibodies with real-valued groundtruth. Given the sequences of the antibody’s heavy and light chains, we need to predict its developability (continuous value). The input consists of a list containing two sequences: the first representing the heavy chain and the second representing the light chain. It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/develop/).
    • SAbDab-Chen (Chen et al., 2020). It consists of 2,409 antibodies with real-valued groundtruth. It is extracted from SAbDab (the structural antibody database) [^3] It is publicly available http://opig.stats.ox.ac.uk/webapps/newSAbDab/SAbDab/. , which is a database containing all the antibody structures available in the PDB (Protein Data Bank), annotated and presented in a consistent fashion (Dunbar et al., 2014). Given the antibody’s heavy chain and light chain sequence, predict its developability (binary label). It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/develop/).
  • CRISPR Repair Outcome Prediction. CRISPR-Cas9 is a gene editing technology that allows for the precise deletion or modification of specific DNA regions within an organism. It operates by utilizing a custom-designed guide RNA that binds to a target site upstream, which results in a double-stranded DNA break facilitated by the Cas9 enzyme. The cell responds by activating DNA repair mechanisms, such as non-homologous end joining, leading to a range of gene insertion or deletion mutations (indels) of varying lengths and frequencies. This task aims to predict the outcomes of these repair processes based on the DNA sequence. Gene editing marks a significant advancement in the treatment of challenging diseases that conventional therapies struggle to address, as demonstrated by the FDA’s recent approval of gene-edited T-cells for the treatment of acute lymphoblastic leukemia. Since many human genetic variants linked to diseases arise from insertions and deletions, accurately predicting gene editing outcomes is essential for ensuring treatment effectiveness and reducing the risk of unintended pathogenic mutations.
    • CRISPR-Leenay (Leenay et al., 2019). The dataset comprises 1,521 DNA sequences (including guide RNA and PAM) with five measured repair outcomes, assessed across various donor populations of primary T cells. It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/CRISPROutcome/).
In this library, we follow the train-validation-test split in PEER benchmark (Xu et al., 2022) and TDC (Huang et al., 2022). Each individual split is reported from Table 2 to 6.
3.2 Cutting-edge Deep Learning Methods
At the core of deep learning lies the artificial neural network, a machine learning technique inspired by the architecture and functionality of the human brain. What distinguishes deep learning from other machine learning approaches is its exceptional capacity to recognize and analyze complex, nonlinear patterns in data, leading to enhanced performance and accuracy. Concretely, we incorporate several cutting-edge neural network architectures into two groups: 1) sequential-based learning and 2) structural-based learning. Detailed model architectures are described as follows:
Sequential based learning
It generally takes a sequence as an input, uses one-hot encoding to pre-encode the input characters. Such learning methods include convolutional neural network, recurrent neural network and transformers.
  • Convolutional Neural Network (CNN) (One-dimensional) captures the local patterns in the data features, commonly used to analyze images and text. (One-dimensional) Convolutional neural network (CNN) takes amino acid sequences as the input. CNN has four layers; the number of filters for the four layers is 32, 64, and 96 respectively. The kernel sizes are 4, 8, and 12, respectively. The convolutional layer is followed by a one-layer MLP (multi-layer perceptron) to produce the prediction, which is a scalar.
  • Recurrent Neural Network (RNN) models sequence data and captures the long-term dependencies in the sequence data. RNN has two well-known variants: long short-term memory networks (LSTMs) (Hochreiter and Schmidhuber, 1996) and gated recurrent units (GRU) (Cho et al., 2014). The difference between GRU and LSTM is that GRU simplifies LSTM by removing the cell state and reducing the number of gates. We use a two-layer bi-directional GRU following three-layer CNN as the neural network architecture. The dimension of the hidden state in GRU is set to 64. ReLU function is applied after each GRU or CNN layer.
  • Transformer (Vaswani et al., 2017) architecture leverages the power of self-attention mechanisms and parallel computation to enhance the neural network’s capability and efficiency in handling sequence data. We use the transformer encoder to represent the amino acid sequence. Two layers of transformer architectures are stacked. The dimension of embedding in the transformer is set to 64. The number of attention heads is set to 4. The ReLU function is applied after each self-attention layer. LayerNorm is applied after MLP layers.
    Structural-based learning
    It generally transforms the input sequence into a valid SMILES string, then transforms the chemical substance into a graph. Then, graph filters are learned toward the input graph signal. Such learning methods are widely called Graph Neural Networks. Recently, graph transformers have shown their power in protein function prediction, and we included them as a part of structural-based learning.
  • Graph Neural Network (GNN) is a neural network architecture designed to process graph-structured data that takes input from nodes and edges, facilitating the flow of information between connected components to capture their interactions. It learns vector representations for both individual graph nodes and the overall graph structure. We consider the following GNN variants:
    • Graph Convolutional Network (GCN) (Kipf and Welling, 2016). GCN is a GNN variant that iteratively updates the node representation by aggregating the information from its neighbors. GCN has three layers, and the node embedding dimension is set to 64. After GCN, all the node embeddings are aggregated with a readout function (Weighted Sum and Max) to get graph-level embedding, followed by a one-layer MLP to get the final prediction. BatchNorm is applied after MLP layers.
    • Graph Attention Network (GAT) (Velickovic et al., 2018). GAT employs an attention mechanism to introduce anisotropy into the neighborhood aggregation function. This network features a multi-headed architecture that enhances its learning capacity. The node embedding dimension is 64. Readout function is the same as the one deployed in GCN model.
    • Message Passing Neural Network (MPNN) (Gilmer et al., 2017). MPNN is a GNN variant that considers passing messages (and modeling interactions) between both edges and nodes based on their neighbors. Edge features are included necessarily compared with GCN and GAT. Readout function is Sum And Max. Node and edge embedding dimension is 64.
    • Neural Fingerprint (NeuralFP) (Duvenaud et al., 2015). NeuralFP uses Graph convolutional network (GCN) (Kipf and Welling, 2016) to learn a neural network-based molecular embedding (also known as molecular neural fingerprint, or NeuralFP) from a large amount of molecule data without labels. The neural fingerprint is essentially a real-valued vector, also known as embedding. Then, the neural fingerprint is fixed and fed into a three-layer MLP to make the prediction. Node embedding dimension is 64. BatchNorm is applied after MLP layers.
    • Attentive Fingerprint (AttentiveFP) (Xiong et al., 2019). AttentiveFP is a variant of graph neural networks that is enhanced by the attention mechanism when evaluating node and edge embedding. The model consists of three AttentiveFP layers with individual readout function: AttentiveFP readout. Node and edge embedding dimension is 64.
  • Graph Transformer (Yun et al., 2019) is a type of neural network architecture designed to process graph-structured data by leveraging self-attention mechanisms. They extend the principles of traditional transformers, enabling them to capture the relationships and interactions between nodes in a graph effectively.
    • Path-Augmented Graph Transformer (PAGTN(Chen et al., 2019). It used augmented path features to capture long-range ($>$1 hop) graph properties. The model consists of 5 PAGTN layers with LeakyReLU activation. Node embedding dimension is 64.
    • Graphormer (Ying et al., 2021). It utilized transformer on graphs with spatial, centrality, and edge encoding. For simplicity and scalability on large graphs, we only deployed one Graphormer layer with ReLU activation. Node embedding dimension is 64. LayerNorm is applied after MLP layers.
Training setup.
For all the models, the maximal training epoch number is set to 100. We employed the Adam optimizer (Kingma and Ba, 2014) for training, with a default learning rate of 0.0001 for sequence-based learning and 0.00001 for structural-based learning. The batch size is equal to 32. More detailed hyper-parameter setups are listed in Table 7 in the appendix.
3.3 Experimental Setup and Implementation Details
Code Base.
This library is an extension of the well-established drug discovery library, DeepPurpose (Huang et al., 2020), building upon its foundational capabilities to offer enhanced features for protein-related tasks. By leveraging the strengths of DeepPurpose, this new library provides additional tools and functionalities tailored specifically for protein science. The library is publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
Hardware Configuration.
All experiments that are mentioned in this paper were trained on a 40GB NVIDIA A40 and a 24GB NVIDIA RTX 3090. The parameters we provide have ensured the scalable training on these two types of GPUs. When running GNNs on protein localization tasks, we observed a large portion of GPU memory occupied irregularly, so we recommend cutting down the size of the number of workers from 8 to 4 or batch size from 32 to 8 or even smaller to potentially avoid GPU out-of-memory (OOM) problems.
Software Configuration.
The library is implemented in Python 3.9, PyTorch 2.3.0, PyTDC 0.4.1 (Huang et al., 2021), DeepPurpose 0.1.5 (Huang et al., 2020), and RDKit 2023.9.6 (Landrum et al., 2006), scikit-learn 1.2.2 (Pedregosa et al., 2011), and DGLlife 0.3.2 (Li et al., 2021a). Besides, wandb is included in DeepProtein so that researchers can observe the visualization of training curves and test results easily.
Figure 2: Performance of selected models including CNN, Transformer, GCN and NeuralFP on SubCellular, PPI Affinity and CRISPR Leenay. We observe that CNN performs significantly better than GCN.
Table 2: Results of protein function prediction. The $\uparrow$ symbol indicates that higher values are better for the corresponding metric. For each method, we employed five different random seeds to perform independent runs, reporting the average results along with their standard deviations. On each task, the best method is bolded, and the second best is underlined. We use “**” to denote the method that achieves statistically better results than all the other methods (through statistical tests).
Model Fluorescence ($\rho$uparrow) Stability ($\rho$uparrow) $\beta$-lactamase ($\rho$uparrow) Solubility (PR-AUC $\uparrow$)
# train/valid/test 21446 / 5362 / 27217 53571 / 2512 / 12851 4158 / 520 / 520 62478 / 6942 / 1999
CNN 0.680 $\pm$ 0.001 0.715 $\pm$ 0.025** 0.721 $\pm$ 0.020** $74.61\pm 0.55$
CNN-RNN $0.678\pm 0.001$ $0.638\pm 0.021$ $0.695\pm 0.012$ $74.10\pm 1.48$
Transformer $0.648\pm 0.001$ $0.442\pm 0.030$ $0.348\pm 0.002$ 78.86 $\pm$ 0.46**
GCN $0.397\pm 0.002$ $0.392\pm 0.027$ $0.417\pm 0.001$ $69.26\pm 0.73$
GAT $0.251\pm 0.002$ $0.165\pm 0.026$ $0.196\pm 0.013$ $62.44\pm 0.01$
NeuralFP $0.413\pm 0.011$ $0.333\pm 0.041$ $0.133\pm 0.020$ $78.74\pm 0.24$
AttentiveFP $0.260\pm 0.003$ $0.254\pm 0.022$ $0.058\pm 0.011$ $78.39\pm 0.19$
MPNN $0.237\pm 0.020$ $0.110\pm 0.021$ $0.068\pm 0.015$ $62.53\pm 0.31$
PAGTN $0.188\pm 0.036$ $0.266\pm 0.016$ $0.092\pm 0.018$ $61.33\pm 0.91$
Graphormer $0.067\pm 0.002$ OOM $0.103\pm 0.018$ OOM
Table 3: Results of protein localization prediction.
Model Subcellular (Acc $\uparrow$) Binary (PR-AUC $\uparrow$)
# train/valid/test 8,945 / 2248 / 2768 5161 / 1727 / 1746
CNN 50.18 $\pm$ 1.21 $90.88\pm 0.31$
CNN-RNN 52.58 $\pm$ 0.11** 91.43 $\pm$ 0.45**
Transformer $42.63\pm 0.68$ $78.38\pm 0.25$
GCN $47.45\pm 0.47$ $83.43\pm 0.10$
GAT $45.14\pm 0.10$ $82.15\pm 0.41$
NeuralFP $45.20\pm 0.49$ $81.14\pm 0.06$
AttentiveFP $42.38\pm 1.25$ $80.58\pm 0.30$
Table 4: Results of Protein-Protein Interaction (PPI).
Model PPI Affinity ($R^{2}$uparrow) Yeast (PR-AUC $\uparrow$) Human PPI (PR-AUC $\uparrow$)
# train/valid/test 2127 / 212 / 343 1668 / 131 / 373 6844 / 277 / 227
CNN $0.493\pm 0.015$ $51.93\pm 0.92$ $70.37\pm 1.22$
CNN-RNN 0.584 $\pm$ 0.026** $53.28\pm 0.85$ $70.45\pm 2.68$
Transformer $0.425\pm 0.021$ $53.79\pm 1.07$ $59.36\pm 4.00$
GCN $0.394\pm 0.006$ 58.98 $\pm$ 0.72 82.21 $\pm$ 1.13
GAT $0.230\pm 0.001$ $53.72\pm 0.39$ $77.63\pm 3.13$
NeuralFP $0.100\pm 0.054$ $57.00\pm 1.51$ $80.11\pm 1.25$
Table 5: Results of epitope and paratope prediction (residue-level classification).
Model IEDB (ROC-AUC $\uparrow$) PDB-Jespersen (ROC-AUC $\uparrow$) SAbDab-Liberis (ROC-AUC $\uparrow$)
# train/valid/test 2211 / 316 / 632 313 / 45 / 89 716 / 102 / 205
CNN $54.03\pm 0.02$ 74.46 $\pm$ 0.21** 90.85 $\pm$ 0.08
CNN-RNN $55.47\pm 0.23$ $70.10\pm 0.97$ 96.75 $\pm$ 0.10**
Transformer 59.79 $\pm$ 0.06** $60.10\pm 0.32$ $64.77\pm 0.11$
Table 6: Results of antibody developability prediction (TAP and SAbDab-Chen) and CRISPR repair outcome prediction (CRISPR-Leenay).
Model TAP (MAE $\downarrow$) SAbDab-Chen (MAE $\downarrow$) CRISPR-Leenay (MAE $\downarrow$)
# train/valid/test 169 / 24 / 48 1686 / 241 / 482 1065 / 152 / 304
CNN $3.217\pm 0.026$ 0.219 $\pm$ 0.006 0.0745 $\pm$ 0.0005
CNN-RNN 0.712 $\pm$ 0.069** $0.226\pm 0.001$ $0.0755\pm 0.0010$
Transformer $3.476\pm 0.004$ $0.238\pm 0.012$ $0.1216\pm 0.0020$
GCN $2.761\pm 0.155$ $0.326\pm 0.015$ $0.1274\pm 0.0019$
GAT $2.675\pm 0.022$ $0.310\pm 0.010$ $0.1232\pm 0.0001$
NeuralFP $3.436\pm 0.015$ $0.253\pm 0.011$ $0.1243\pm 0.0001$
3.4 Results & Analysis
For each method, we used five different random seeds to conduct independent runs and reported the average results and their standard deviations. The results of protein function prediction are reported in Table 2.
Statistical Test. We also conduct statistical tests to confirm the superiority of the best-performed method compared with the second-best baseline method. The hypothesis is that the accuracies of the best method are the same as those of the baseline method. Student’s T-test is used with significance level alpha as 1% to calculate the p-values. When the p-values are below the 0.05 threshold, we reject the hypothesis and accept the alternative hypothesis, i.e., the best method is statistically significant compared with the second-best method. We use “**” to denote the method that achieves statistically better results than all the other methods (pass statistical tests).
Key Observations.
We summarize the following key observations as takeaways.
  • Sequence-based neural architectures are powerful compared with graph-based neural architectures. Sequence-based neural architectures, such as CNN, RNN, and transformer, obtain superior performance in most protein sequence learning tasks. Specifically, in 12 out of all the 15 tasks across various protein sequence learning tasks, the sequence-based model (CNN, RNN, and transformer) takes the top-2 position.
  • Among all the 12 GNN-solvable tasks (except residue-level classification), graph neural networks (GNN) obtain the best accuracy among all the methods only in two protein-protein interaction (PPI) tasks, including Yeast and Human PPI.
  • Among all the graph neural networks (GNNs) across the whole 12 GNN-solvable tasks (except residue-level classification), the earliest variant, GCN (Kipf and Welling, 2016), achieves the best performance in 9 tasks.
  • Stability. From the learning curve (Figure 7-10), we find that GNN’s training curve is not stable. In contrast, the sequence-based models, including CNN, RNN, and transformer, converge more stably from the learning curve. This can be observed from Figure 7-9. On the contrary, training is more stable, fast and accurate for GAT when it comes to TAP dataset.
  • Computational complexity. The runtime and memory costs are reported in Figure 35 and 6. We find that GNN-based models are typically computationally inefficient. The key reason behind this is that GNN utilizes molecular graph as the feature, where each atom corresponds to a node and each chemical bond corresponds to an edge. While another model, such as CNN, RNN, and transformer, uses amino acid sequences as the input feature.
Figure 3: Empirical computational complexity on CRISPR and SAbDab Chen measured by runtime and memory.
4 Conclusion
In this paper, we have developed DeepProtein, which marks a significant advancement in the application of deep learning to protein science, providing researchers with a powerful and flexible tool to tackle various protein-related tasks. By integrating multiple state-of-the-art neural network architectures and offering a comprehensive benchmarking suite, DeepProtein empowers users to explore and optimize their models effectively. The detailed documentation and tutorials further enhance accessibility, promoting widespread adoption and reproducibility in research. As the field of proteomics continues to evolve, DeepProtein stands to contribute substantially to our understanding of protein functions, localization, and interactions, ultimately driving forward discoveries that can impact biotechnology and medicine.
Appendix A Hyperparameter Settings
In table 7, we have listed a common settings of hyperparameter used in this library. In terms of learning rate (lr), a higher learning rate which is equal to 0.0001 for graph neural networks would lead to failure in training. For Subcellular and its binary version, a training epoch of 60 is enough for convergence. For small-scale protein datasets such as IEDB [Yi et al., 2018], PDB-Jespersen, and SAbDab-Liberis, a larger learning rate of 0.001 also leads to convergence and the same performance when using CNN, CNN-RNN, and Transformer. For TAP, SAbDab-Chen and CRISPR-Leenay, larger learning rate of 0.0001 is suggested when training graph neural networks.
Table 7: Default Model Configurations for Protein Sequence Learning.
Model lr dropout activation # heads # layers hidden dim pooling batch size # epochs norm
CNN $10^{-4}$ 0.1 ReLU - 3 256 MaxPool1d 32 100 -
CNN-GRU $10^{-4}$ 0.1 ReLU - 2 64 - 32 100 -
Transformer $10^{-4}$ 0.1 ReLU 4 2 64 - 32 100 LayerNorm
GCN $10^{-5}$ 0.1 ReLU - 3 64 Weighted Sum + Max 32 100 BatchNorm
GAT $10^{-5}$ 0.1 ReLU - 3 64 Weighted Sum + Max 32 100 -
NeuralFP $10^{-5}$ 0.1 ReLU - 3 64 Sum + Max 32 100 BatchNorm
AttentiveFP $10^{-5}$ 0.1 ReLU - 3 64 AttentiveFPReadout 32 100 -
MPNN $10^{-5}$ 0.1 ReLU - 6 64 Sum + Max 32 100 -
PAGTN $10^{-5}$ 0.1 LeakyReLU - 5 64 Weighted Sum + Max 32 100 -
Graphormer $10^{-5}$ 0.1 ReLU 8 1 64 MaxPooling 32 100 LayerNorm
Appendix B More results of performance on protein sequence learning
In this section, we show three more model performance on Fluorescence, Beta-lactamase and Human PPI dataset.
Figure 4: Performance of selected models including CNN, Transformer, GCN and NeuralFP on Fluorescence, Beta-lactamase and Human PPI. We observe that CNN performs significantly better than GCN on protein function prediction tasks. However, GCN dominates on large-scale PPI datasets such as the Human PPI dataset.
Appendix C More results of empirical memory and runtime complexity
In this section, we show four more empirical complexity results on Fluorescence, Beta-lactamase, SubCellular Binary and Yeast PPI. In Figure 5, the same pattern is found as Figure 3: Transformer took up a large amount of memory to train while the training speed is reasonably fast. Training GNN is slow since the DGL framework computes the message passing on the CPU. CNNs are both fast and memory efficient to train. A different case is found in Figure 6 where for SubCellular Binary and Yeast PPI, GCN is the slowest to train with occupying a large amount of GPU memory space.
Figure 5: Empirical computational complexity on CRISPR and SAbDab Liberis measured by runtime and memory.
Figure 6: Empirical computational complexity on SubCellular Binary and Yeast PPI measured by runtime and memory.
Appendix D Training Curves of DeepProtein
In this section, we plot the metrics (Accuracy or R2) and training loss for SubCellular, PPI Affinity, SAbDab-Chen and TAP.
Figure 7: SubCellular
Figure 8: PPI Affinity
Figure 9: SAbDab_Chen
Figure 10: TAP
Appendix E Few Lines of Codes
Protein Function Prediction
We show an example of CNN on Beta_lactamase with DeepProtein. More case studies can be found in the GitHub repository of DeepProtein. We annotate the code to make researchers better understand the function of code base, which is simpler and clearer than most of the protein learning libraries available.
$>$ import os, sys, argparse, torch, wandb
$>$ from DeepProtein.dataset import *
$>$ import DeepProtein.utils as utils
$>$ import DeepProtein.ProteinPred as models
$>$ path = os.getcwd()
$>$ train_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’train’)
$>$ valid_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’valid’)
$>$ test_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’test’)
$>$ train_protein_processed,train_target,train_protein_idx = collate_fn(train_beta)
$>$ valid_protein_processed,valid_target,valid_protein_idx = collate_fn(valid_beta)
$>$ test_protein_processed,test_target,test_protein_idx = collate_fn(test_beta)
$>$ target_encoding = ’CNN’
$>$ train, _, _ = utils.data_process(X_target=train_protein_processed, y=train_target, target_encoding=target_encoding, split_method=’random’, frac=[0.99998, 1e-5, 1e-5])
$>$ _, val, _ = utils.data_process(X_target=valid_protein_processed, y=valid_target, target_encoding=target_encoding, split_method=’random’, frac=[1e-5, 0.99998, 1e-5])
$>$ _, _, test = utils.data_process(X_target=test_protein_processed, y=test_target, target_encoding=target_encoding,split_method=’random’, frac=[1e-5, 1e-5, 0.99998])
$>$ config = generate_config(target_encoding=target_encoding, \
cls_hidden_dims=[1024, 1024], \
train_epoch=20, \
LR=0.0001, \
batch_size=32, \
$>$ config[’multi’] = False
$>$ torch.manual_seed(args.seed)
$>$ model = models.model_initialize(**config)
$>$ model.train(train, val, test, compute_pos_enc = False)
If we use argparse in Python, this can be easily done in one line, where
python beta.py --target_encoding CNN
--seed 7
--wandb_proj DeepPurposePP
--lr 0.0001
--epochs 100
Note that if you want to observe the result offline without wandb, then set
wandb.init(mode=’offline’, name = "your_project")
inside the training file.
Appendix F Tutorial
Demos can be found under DEMOS folder in DeepProtein.

### v2

@@
 Abstract
-In recent years, deep learning has revolutionized the field of protein science, enabling advancements in predicting protein properties, structural folding and interactions.
-This paper presents DeepProtein, a comprehensive and user-friendly deep learning library specifically designed for protein-related tasks.
-DeepProtein integrates a couple of state-of-the-art neural network architectures, which include convolutional neural network (CNN), recurrent neural network (RNN), transformer, graph neural network (GNN), and graph transformer (GT).
-It provides user-friendly interfaces, facilitating domain researchers in applying deep learning techniques to protein data.
-Also, we curate a benchmark that evaluates these neural architectures on a variety of protein tasks, including protein function prediction, protein localization prediction, and protein-protein interaction prediction, showcasing its superior performance and scalability.
-Additionally, we provide detailed documentation and tutorials to promote accessibility and encourage reproducible research.
-This library is extended from a well-known drug discovery library, DeepPurpose and publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
+Deep learning has deeply influenced protein science, enabling breakthroughs in predicting protein properties, higher-order structures, and molecular interactions.
+This paper introduces DeepProtein, a comprehensive and user-friendly deep learning library tailored for protein-related tasks.
+It enables researchers to seamlessly address protein data with cutting-edge deep learning models.
+To assess model performance, we establish a benchmark evaluating different deep learning architectures across multiple protein-related tasks, including protein function prediction, subcellular localization prediction, protein-protein interaction prediction, and protein structure prediction.
+Furthermore, we introduce DeepProt-T5, a series of fine-tuned Prot-T5-based models that achieve state-of-the-art performance on four benchmark tasks, while demonstrating competitive results on six of others.
+Comprehensive documentation and tutorials are available which could ensure accessibility and support reproducibility.
+Built upon the widely used drug discovery library DeepPurpose, DeepProtein is publicly available at https://github.com/jiaqingxie/DeepProtein.
 1 Introduction
-Understanding the representation of proteomics is vital in developing traditional biological and medical progress (Wu et al., 2022b; Fu et al., 2024), multi-omics genomics (Wu et al., 2022a; Chen et al., 2021), and curing human diseases (Chen et al., 2024c, b).
-Being the working house of the cell, it provides lots of functions that support human’s daily life, such as catalyzing biochemical reactions that happen in the body as a role of enzymes and providing helpful immune responses against detrimental substance that acts as immunoglobulin (Wu et al., 2024).
-Under the necessity of analyzing those useful proteins, several related protein databases are available to researchers (Berman et al., 2000; Bairoch and Apweiler, 2000; Consortium, 2015; Pontén et al., 2008).
+Understanding the representation of proteomics is vital in developing traditional biological and medical progress (Wu et al., 2022b; Fu et al., 2024), multi-omics genomics (Wu et al., 2022a; Chen et al., 2021), and curing human diseases (Chen et al., 2024b; c).
+Being the working house of the cell, it provides many functions that support human daily life, such as catalyzing biochemical reactions that occur in the body as a role of enzymes and providing helpful immune responses against harmful substances that act as immunoglobulin (Wu et al., 2024).
+Under the necessity of analyzing those useful proteins, several related protein databases are available to researchers (Berman et al., 2000; Bairoch & Apweiler, 2000; Consortium, 2015; Pontén et al., 2008).
 Apart from the 2D database, some recent 3D Protein Database used AlphaFold 2.0 (Jumper et al., 2021) is important to better assist in learning those representations in 3d-dimensional space.
 The success of AlphaFold 2.0 has sparked a significant increase in interest in using machine learning techniques for protein learning tasks, of which the goal is to improve our understanding of proteins’ biochemical mechanisms.
-Deep learning has shown its power in protein learning tasks, including protein-protein interaction (Gainza et al., 2020), protein folding (Jumper et al., 2021; Lu, 2022; Panou and Reczko, 2020; Chen et al., 2016), protein-ligand interaction (Li et al., 2021b; Xia et al., 2023), and protein function and property prediction (Gligorijević et al., 2021; Sevgen et al., 2023).
-Convolutional neural network (CNN) (Shanehsazzadeh et al., 2020) and TAPE Transformer (Rao et al., 2019) on proteins are models with regard to protein sequence-to-sequence (Seq2Seq) learning, including the pretrained transformer such as ProtBert (Brandes et al., 2022).
-With the help of 3D information intrinsics, graph neural networks (GNN) are applied, especially some pretrained GNN models (Jing et al., 2020; Zhang et al., 2022), and graph transformers (Yuan et al., 2022; Gu et al., 2023).
-While transformer claims to be the state-of-the-art in the previous benchmark (Xu et al., 2022), the analysis of potential powerful deep learning models such as GNN models and graph transformer models are neglected, which prompts us to integrate those methods in our benchmark.
-Also, pretaining strategies are well-known for its ability in utilizing unlabeled data (Lu et al., 2024; Yue et al., 2024).
-Challenges.
-Previous benchmarks related to molecular learning have offered valuable insights regarding their respective libraries and implementation interfaces.
-DeepPurpose [^1] https://github.com/kexinhuang12345/DeepPurpose (Huang et al., 2020) has provided an interface that implements the task with a majority of drug discovery tasks, which only has protein-protein interaction and protein function prediction implemented.
-Datasets on proteins are lacking as well.
-TorchProtein [^2] https://github.com/DeepGraphLearning/PEER_Benchmark  (Xu et al., 2022), also named as PEER, implemented most of the tasks in the protein field.
-In terms of models, the focus has largely been on sequential learning methods: Convolutional Neural Networks (CNNs), Transformers, and ESM architectures.
-This suggests that there are still many structural methods available for consideration.
-Furthermore, PEER’s interface is not user-friendly without prior domain knowledge in graphs and biochemistry.
-This situation presents an opportunity to improve the existing interface with regard to simplicity and comprehensibility.
-Solutions.
-To address these challenges, in this paper, we propose DeepProtein, which aims to benchmark mainstream and cutting-edge deep learning models on a wide range of AI-solvable protein sequence learning tasks.
-We investigate the performance of various deep learning models on a wide range of protein sequence learning tasks.
-We analyze each method’s advantages and disadvantages when performing each task (working as the explainer for each task).
-We have provided user-friendly and well-wrapped interfaces to facilitate domain experts’ research.
-Contribution. The major contributions of our paper are summarized as
-  • Comprehensive Benchmarking: We curate a benchmark to evaluate the performance of 8 neural network architectures (including CNNs, RNNs, transformers, and various graph neural networks) on these 7 essential protein learning tasks (including protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, antibody paratope prediction, CRISPR repair outcome prediction, and antibody developability prediciton), demonstrating superior performance and scalability.
-  • User-friendly Library: We develop DeepProtein, a specialized deep learning library that integrates these neural network architectures (CNNs, RNNs, transformers, and graph neural networks) on these 7 protein-related tasks such as protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, and antibody developability prediciton.
-For each task, our library enables one command line to run every method.
-  • Enhanced Accessibility: We provide extensive documentation and tutorials to facilitate user engagement and promote reproducible research, building upon the foundation of the well-known drug discovery library, DeepPurpose (Huang et al., 2020).
-    Figure 1: DeepProtein framework.
-2 Related Works
-Benchmarks and libraries are crucial in AI-based therapeutic science, e.g., multi-omics data (Lu, 2018), protein learning (Xu et al., 2022), small-molecule drug discovery (Gao et al., 2022; Zheng et al., 2024; Xu et al., 2024), and drug development (clinical trial) (Chen et al., 2024a; Wang et al., 2024).
-They provide standardized metrics for evaluating the performance of various algorithms and models.
-These benchmarks enable researchers to compare different approaches systematically, ensuring reproducibility and reliability of results.
-In this section, we briefly discuss the benchmark studies in this area.
-Proteins are vital in drug discovery because they often serve as the primary targets for therapeutic agents, influencing disease mechanisms and biological pathways.
-Additionally, proteins play key roles in various cellular processes, making them essential for identifying potential drug candidates and biomarkers in the drug development pipeline.
-A couple of protein learning benchmarks are developed, including PEER (Xu et al., 2022), DeepPurpose (Huang et al., 2020), FLIP (Dallago et al., 2021), TAPE (Rao et al., 2019).
-Table 1 compares DeepProtein with existing AI-based protein learning benchmarks.
-We extend the scope of existing protein learning benchmarks by incorporating more protein learning datasets, more cutting-edge deep learning models, and enhancing user-friendliness.
+Deep learning has revolutionized protein science, driving significant advancements in various protein-related tasks.
+These include protein-protein interactions (Gainza et al., 2020), protein folding (Jumper et al., 2021; Lu, 2022; Panou & Reczko, 2020; Chen et al., 2016), protein-ligand interactions (Li et al., 2021b; Xia et al., 2023), and protein function and property prediction (Gligorijević et al., 2021; Sevgen et al., 2023).
+The development of deep neural architectures has played a crucial role in these tasks, with approaches leveraging both sequence-based and structure-based models.
+Sequence-based models, such as convolutional neural networks (CNNs) (Shanehsazzadeh et al., 2020) and transformers, have shown strong performance in protein learning tasks.
+The TAPE Transformer (Rao et al., 2019) and pre-trained transformer models such as ProtBERT (Brandes et al., 2022) have demonstrated the effectiveness of self-supervised learning in capturing protein sequence representations.
+Beyond sequence-based methods, structure-based deep learning has gained traction with graph neural networks (GNNs), which leverage 3D structural information to enhance structural property (Jing et al., 2020; Zhang et al., 2022).
+Recently, graph transformers have emerged as a powerful alternative, combining the advantages of transformers (global attention) and message-passing neural networks (sparse attention) to model protein structures more effectively (Yuan et al., 2022; Gu et al., 2023).
+While transformers have been considered state-of-the-art in previous benchmarks (Xu et al., 2022), comprehensive comparisons between CNN, transformer, GNN, and other advanced architectures remain under-explored.
+This gap motivates us to systematically integrate and evaluate these methods in our benchmark.
+Furthermore, pretraining strategies have been prevailing in protein science, which have utilized the large-scale unlabeled protein data to improve downstream performance (Lu et al., 2024; Yue et al., 2024).
+With the advent of large foundation models, protein properties can now be inferred through prompt engineering, such as BioMistral (Labrak et al., 2024), BioT5/BioT5+ (Pei et al., 2023; 2024), and ChemLLM (Zhang et al., 2024b).
+Both advancements in molecule pretraining and question-answering language models in molecules brought more possibilities in the field of protein engineering.
 […]
@@
 […]
-user-friendly $\times$ $\times$ $\times$
+Fold $\times$ $\times$ $\times$ $\times$
+Secondary Structure $\times$ $\times$ $\times$
+Challenges.
+Previous benchmarks related to molecular learning have offered valuable insights regarding their respective libraries and implementation interfaces.
+DeepPurpose [^1] https://github.com/kexinhuang12345/DeepPurpose (Huang et al., 2020) has provided an interface that implements the task with a majority of drug discovery tasks, which only has protein-protein interaction and protein function prediction implemented.
+Datasets on proteins are lacking as well.
+TorchProtein [^2] https://github.com/DeepGraphLearning/PEER_Benchmark  (Xu et al., 2022), also named as PEER, implemented most of the tasks in the protein field.
+In terms of models, the focus has largely been on sequence-based methods: Convolutional Neural Networks (CNNs), Transformers, and ESM architectures.
+This suggests that there are still many structure-based methods (GNN) or pre-trained protein language models available (such as ProtBert or Prot-T5) for consideration.
+Furthermore, PEER’s interface is not user-friendly without prior domain knowledge in graphs and biochemistry.
+This presents an opportunity to improve the existing interface regarding simplicity and comprehensibility.
+Solutions.
+To address these challenges, in this paper, we propose DeepProtein, which aims to benchmark mainstream and cutting-edge deep learning models on a wide range of AI-solvable protein sequence learning tasks.
+We investigate the performance of various deep learning models on a wide range of protein sequence learning tasks.
+We analyze each method’s advantages and disadvantages when performing each task (working as the explainer for each task).
+We have provided user-friendly and well-wrapped interfaces to facilitate domain experts’ research.
+Contribution. Our key contributions are summarized as:
+  • Comprehensive Benchmarking: We curate a benchmark to evaluate the performance of eight coarse grained deep learning architectures, including CNNs, CNN-RNNs, RNNs, transformers, graph neural networks, graph transformers, pre-trained protein language models, and large language models.
+This benchmark covers eight protein learning tasks, including protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, antibody paratope prediction, CRISPR repair outcome prediction, antibody developability prediction, and protein structure prediction.
+Our benchmark demonstrates the strength, scalability and limitation of the mentioned approach, respectively.
+  • User-friendly Library: We develop DeepProtein, a specialized deep learning library that integrates these neural network architectures for protein-related tasks.
+DeepProtein offers a simple, command-line interface for running models on all supported tasks, making it accessible to researchers with minimal deep learning expertise.
+  • Enhanced Accessibility: We provide comprehensive documentation, tutorials, and pre-configured pipelines.
+Inherited from DeepPurpose (Huang et al., 2020), our library ensures seamless integration with existing protein frameworks or personalized protein databases, and enables reproducibility.
+  • Fine-Tuned Models – DeepProt-T5: We have released our fine-tuned Prot-T5-XL models for each task, which is available on HuggingFace.
+The model family is called DeepProt-T5.
+These models achieve either state-of-the-art or competitive performance across our DeepProtein benchmark, so there is no need for the redundant retraining process, making model deployment much more efficient and convenient.
+2 Related Works
+Benchmarks and libraries are crucial in AI-based therapeutic science, e.g., multi-omics data (Lu, 2018), protein learning (Xu et al., 2022), small-molecule drug discovery (Gao et al., 2022; Zheng et al., 2024; Xu et al., 2024), and drug development (clinical trial) (Chen et al., 2024a; Wang et al., 2024).
+They provide standardized metrics for evaluating the performance of various algorithms and models.
+These benchmarks enable researchers to compare different approaches systematically, ensuring reproducibility and reliability of results.
+In this section, we briefly discuss the benchmark studies in this area.
+Proteins are vital in drug discovery because they often serve as the primary targets for therapeutic agents, influencing disease mechanisms and biological pathways.
+Additionally, proteins play key roles in various cellular processes, making them essential for identifying potential drug candidates and biomarkers in the drug development pipeline.
+A couple of protein learning benchmarks are developed, including PEER (Xu et al., 2022), DeepPurpose (Huang et al., 2020), FLIP (Dallago et al., 2021), TAPE (Rao et al., 2019).
+Table 1 compares DeepProtein with existing AI-based protein learning benchmarks.
+We extend the scope of existing protein learning benchmarks by incorporating more protein learning datasets, more cutting-edge deep learning models, and enhancing user-friendliness.
+Figure 1: DeepProtein framework.
+Part 1.
+A.
+DeepProtein is mainly selected from TorchDrug and Therapeutics Data Commons (TDC), where only protein tasks are considered, and all drug-related tasks are excluded, such as drug-target interactions.
+Specifically, DeepPurpose has established such a pipeline in their library.
+B.
+Both sequence-based and structure-based methods are included in DeepProtein.
+For some graph neural networks, we utilized edge featurizers to generate additional edge information since inputs are 2-dimensional.
+Protein language models, large language models, and our pre-trained T5 (DeepProt-T5) are discussed in Figure 2.
+C.
+Task types are: protein function prediction, subcellular localization prediction, protein-protein interaction prediction, and protein structure prediction.
+They can be classified as either a 1 (protein)-to-1 (aim) problem or a 2 (proteins)-to-1 (aim) problem, which meets the researchers’ needs.
+D.
+An earlier version of DeepProtein could be executed within 20 lines of code.
+The newest version of DeepProtein could be executed within 10 lines of code, where we further wrapped the data processing steps.
+Researchers can also provide their own data to either train or perform inference with the help of DeepProtein.
+E.
+In this paper, we provide comprehensive results, including the performance of each model on corresponding tasks, the differences among sequence-based models, structure-based models, and pre-trained protein language models, and the computation resources, including time-stamps and GPU memory assumptions.
+As DeepProtein supports wandb, we also provide two wandb repositories that record the results of all experiments, which are https://wandb.ai/jiaqing/DeepProtein?nw=nwuserjiaqing and https://wandb.ai/jiaqing/DeepPurposePP.
+Tables and figures are presented later in this paper.
 3 DeepProtein Library and Benchmark
@@
 PPI prediction seeks to forecast the activity of these interactions based on the amino acid sequences of paired proteins.
-    • PPI Affinity (Moal and Fernández-Recio, 2012). It consists of 2,682 protein-protein pairs with real-valued groundtruth.
+    • PPI Affinity (Moal & Fernández-Recio, 2012). It consists of 2,682 protein-protein pairs with real-valued groundtruth.
     • Yeast (Guo et al., 2008). The dataset comprises 2,172 protein-protein pairs with binary labels.
@@
 This is also known as residue-level classification.
-    • SAbDab-Liberis (Liberis et al., 2018) is curated from SAbDab (Dunbar et al., 2014).
-It consists of 1,023 antibody chain sequences; each antibody contains both heavy and light chain sequences.
-It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/paratope/#sabdab-liberis-et-al).
   • Antibody Developability Prediction.
@@
 It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/CRISPROutcome/).
-In this library, we follow the train-validation-test split in PEER benchmark (Xu et al., 2022) and TDC (Huang et al., 2022). Each individual split is reported from Table 2 to 6.
+  • Protein Structure Prediction.
+Protein structure prediction (PSP) is a fundamental problem in computational biology, aiming to determine the three-dimensional structure of a protein from its amino acid sequence.
+Specifically, in our benchmark, the task is to predict the family of a folding or secondary structure family that it belongs to.
+Since a protein’s structure dictates its function, accurate prediction is crucial for understanding the topology of the protein.
+PSP can be broadly divided into global topology prediction and local structural prediction, which include tasks such as fold classification and secondary structure prediction:
+    • Fold (Hou et al., 2018).
+This task involves predicting the global structural topology of a protein, categorizing it into one of the predefined fold classes (within 0, 1, …, 1194).
+During inference, we predict the class of each protein’s superfamily.
+It contains 13766 samples overall.
+    • Secondary Structure (Klausen et al., 2019).
+This task focuses on predicting the local structural elements (coil, strand, helix) of each residue in a protein sequence.
+It serves as an intermediate step for more complex structure prediction tasks and is useful in applications such as functional analysis and multiple sequence alignment.
+This is a residue-level 3-class classification problem where the number of samples is equal to 11361.
+In this library, we follow the train-validation-test split in PEER benchmark (Xu et al., 2022) and TDC (Huang et al., 2022). Each individual split is reported from Table 2 to 7.
+Figure 2: DeepProtein framework, and DeepProt-T5.
+Part 2.
+A.
+For pre-trained protein language models, we directly use the initial string as the input instead of the transformed SMILES , since most of the protein language models have learned such representations.
+Strings are tokenized and carried as inputs to the model.
+We fine-tuned the models on the downstream tasks with the fixed embeddings, which means that after feature extraction, the upstream model parameters will not be trained any more.
+B.
+For large language models, such as ChemLLM and LlaSMol, we did not fine-tune them on the downstream tasks due to limited GPU resources.
+Instead we directly performed inference on downstream tasks where appropriate prompt templates should be carefully designed here.
+C.
+We extend the family of Prot-T5 to our DeepProt-T5, where the upstream embeddings are dynamic, by using the huggingface trainer.
+All fine-tuned models could be found here:https://huggingface.co/jiaxie.
 3.2 Cutting-edge Deep Learning Methods
 At the core of deep learning lies the artificial neural network, a machine learning technique inspired by the architecture and functionality of the human brain.
-What distinguishes deep learning from other machine learning approaches is its exceptional capacity to recognize and analyze complex, nonlinear patterns in data, leading to enhanced performance and accuracy.
+What distinguishes deep learning from other machine learning approaches is its exceptional ability to recognize and analyze complex, non-linear patterns in data, leading to enhanced performance and accuracy.
 Concretely, we incorporate several cutting-edge neural network architectures into two groups: 1) sequential-based learning and 2) structural-based learning.
 Detailed model architectures are described as follows:
-Sequential based learning
-It generally takes a sequence as an input, uses one-hot encoding to pre-encode the input characters.
-Such learning methods include convolutional neural network, recurrent neural network and transformers.
+Sequential based learning It generally takes a sequence as an input and uses one-hot encoding to pre-encode the input characters.
+Such learning methods include convolutional neural networks, recurrent neural networks, and transformers.
   • Convolutional Neural Network (CNN) (One-dimensional) captures the local patterns in the data features, commonly used to analyze images and text.
 (One-dimensional) Convolutional neural network (CNN) takes amino acid sequences as the input.
-CNN has four layers; the number of filters for the four layers is 32, 64, and 96 respectively.
+CNN has four layers; the number of filters for the four layers is 32, 64, and 96, respectively.
 The kernel sizes are 4, 8, and 12, respectively.
-The convolutional layer is followed by a one-layer MLP (multi-layer perceptron) to produce the prediction, which is a scalar.
+The convolutional layer is followed by a one-layer MLP (multi-layer perceptron) to predict as a scalar.
   • Recurrent Neural Network (RNN) models sequence data and captures the long-term dependencies in the sequence data.
-RNN has two well-known variants: long short-term memory networks (LSTMs) (Hochreiter and Schmidhuber, 1996) and gated recurrent units (GRU) (Cho et al., 2014).
+RNN has two well-known variants: long short-term memory networks (LSTMs) (Hochreiter & Schmidhuber, 1996) and gated recurrent units (GRU) (Cho et al., 2014).
 The difference between GRU and LSTM is that GRU simplifies LSTM by removing the cell state and reducing the number of gates.
@@
 LayerNorm is applied after MLP layers.
-    Structural-based learning
-    It generally transforms the input sequence into a valid SMILES string, then transforms the chemical substance into a graph.
+    Structural-based learning It generally transforms the input sequence into a valid SMILES string, then transforms the chemical substance into a graph.
 Then, graph filters are learned toward the input graph signal.
@@
 We consider the following GNN variants:
-    • Graph Convolutional Network (GCN) (Kipf and Welling, 2016).
+    • Graph Convolutional Network (GCN) (Kipf & Welling, 2016).
 GCN is a GNN variant that iteratively updates the node representation by aggregating the information from its neighbors.
@@
     • Neural Fingerprint (NeuralFP) (Duvenaud et al., 2015).
-NeuralFP uses Graph convolutional network (GCN) (Kipf and Welling, 2016) to learn a neural network-based molecular embedding (also known as molecular neural fingerprint, or NeuralFP) from a large amount of molecule data without labels.
+NeuralFP uses Graph convolutional network (GCN) (Kipf & Welling, 2016) to learn a neural network-based molecular embedding (also known as molecular neural fingerprint, or NeuralFP) from a large amount of molecule data without labels.
 The neural fingerprint is essentially a real-valued vector, also known as embedding.
@@
 LayerNorm is applied after MLP layers.
+  • Foundation Model.
+A foundation model is a large-scale, pre-trained machine learning model trained on extensive and diverse datasets, typically using self-supervised or unsupervised learning techniques.
+These models learn generalizable features and patterns from data, allowing them to perform various downstream tasks with minimal task-specific fine-tuning.
+    • ESM.
+The Evolutionary Scale Modeling (ESM) utilizes large-scale pretraining on vast protein sequence data to capture evolutionary relationships and functional patterns within proteins (Lin et al., 2023; Rives et al., 2021).
+It benefits from Masked Language Modeling (MLM) and the transformer architecture.
+In this paper, we consider two ESM variants with different model sizes: ESM-1b and ESM-2-650M.
+The latter incorporates Rotary Position Embedding (RoPE) within the ESM-1 framework.
+We evaluate both models, where the embedding size is 1280.
+    • Prot-T5-XL.
+First introduced in ProtTrans (Elnaggar et al., 2021), Prot-T5-XL-UniRef50 is based on the T5-3B model and was pre-trained on a large corpus of protein sequences using a self-supervised approach.
+A key difference from the original T5 model is the denoising objective: while the original T5-3B model used a span denoising objective, this model employs a BART-like MLM denoising objective.
+The masking probability follows the original T5 training, randomly masking 15% of the amino acids in the input.
+The embedding dimension is 1024.
+    • ProtBert.
+In addition to Prot-T5, ProtTrans includes a model pre-trained on BERT.
+Bidirectional Encoder Representations from Transformers (BERT) is a transformer-based neural network architecture pre-trained on unlabeled sequence data (Devlin et al., 2019).
+A key difference between ProtBert and the original BERT is how sequences are treated as separate documents, eliminating the need for next sentence prediction.
+The masking strategy follows the original BERT training, where 15% of the amino acids in the input are randomly masked.
+The embedding dimension is 1024.
+  • Large Language Model.
+In this paper, we distinguish foundation models as a class of pre-trained protein language models, whereas we define large language models (LLMs) as decoder-only models designed to generate sequential responses regarding the properties of one or multiple proteins given their sequences and a specific prompt.
+Due to computational constraints, fine-tuning 7B-scale models is resource-intensive; therefore, we focus on evaluating their performance instead.
+We consider two Protein LLMs in our study: ChemLLM-7B and LlaSMol-Mistral-7B.
+Their generalization ability to protein-related tasks has not been explored before.
+Additionally we provide the chat and prompt template in the appendix.
+    • ChemLLM-7B.
+The backbone of ChemLLM-7B (Zhang et al., 2024a) is the InternLM2-Base-7B model.
+It was initially trained on the Multi-Corpus-1.7M dataset on Hugging Face and later fine-tuned using instruction-tuning methods on ChemData (7M) and Multi-Corpus (1.7M).
+ChemLLM-7B has demonstrated superior performance over GPT-4 in retrosynthesis and temperature prediction tasks.
+The model provides a set of predefined instruction-following templates, which are used in our study and detailed in the appendix.
+    • LlaSMol-Mistral-7B.
+The backbone of LlaSMol-Mistral-7B (Yu et al., 2024) is Mistral-7B.
+It was fine-tuned using SMolInstruct, a large-scale, high-quality dataset designed for instruction tuning.
+SMolInstruct comprises 14 selected chemistry-related tasks and over three million samples, providing a solid foundation for training and evaluating LLMs in the field of chemistry.
+Specifically, in our experiments, we wrap the input SMILES sequences with $\rangle\langle\text{/PROTEIN}\rangle$PROTEIN ⟩ token pairs to adapt them for protein-related tasks.
+The template is detailed in the appendix.
+  • DeepProt-T5.
+We have trained Prot-T5-XL on our benchmark datasets individually (Figure 2).
+Different from the fixed embeedings, dynamic embeddings enabled us to finetuned the upstream architectures while also maintain a good predictive power for downstream tasks.
+This leads to a series of DeepProt-T5 models.
 Training setup.
-For all the models, the maximal training epoch number is set to 100.
-We employed the Adam optimizer (Kingma and Ba, 2014) for training, with a default learning rate of 0.0001 for sequence-based learning and 0.00001 for structural-based learning.
+For all models, the maximal training epoch number is set to 100.
+We used Adam optimizer (Kingma & Ba, 2014) for training, with a default learning rate of 0.0001 for sequence-based learning and 0.00001 for structural-based learning.
 The batch size is equal to 32.
 More detailed hyper-parameter setups are listed in Table 10 in the appendix.
+For DeepProt-T5 models, fine-tuning has used a more generalized learning rate of 0.00002 for all models, batch size is equal to 10 to avoid memory errors.
+Figure 3: Results of two metrics for selected deep learning methods for DeepProtein Benchmark.
+For regression task, metrics are Spearman (Pearson) Coefficient and Mean Absolute Error (MAE).
+For the binary classification task, metrics are ROC (or PR-AUC) and averaged macro F1.
+For multi classification task, metrics are the accuracy and averaged macro F1.
+Our DeepProt-T5 are competitive among deep learning methods included in our benchmark, and have improved original Prot-T5 models on six tasks: Beta-lactamase, Solubility, SubCellular, PPI_Affinity, CRISPR and Fold.
 3.3 Experimental Setup and Implementation Details
@@
 By leveraging the strengths of DeepPurpose, this new library provides additional tools and functionalities tailored specifically for protein science.
-The library is publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
+The library is publicly available at https://github.com/jiaqingxie/DeepProtein/.
 Hardware Configuration.
 All experiments that are mentioned in this paper were trained on a 40GB NVIDIA A40 and a 24GB NVIDIA RTX 3090.
+For DeepProt-T5 fine-tuning, two 24GB NVIDIA RTX 3090 were used.
 The parameters we provide have ensured the scalable training on these two types of GPUs.
@@
 Besides, wandb is included in DeepProtein so that researchers can observe the visualization of training curves and test results easily.
-Figure 2: Performance of selected models including CNN, Transformer, GCN and NeuralFP on SubCellular, PPI Affinity and CRISPR Leenay. We observe that CNN performs significantly better than GCN.
+More details about environment setup could be found in the GitHub.
 Table 2: Results of protein function prediction.
@@
 For each method, we employed five different random seeds to perform independent runs, reporting the average results along with their standard deviations.
-On each task, the best method is bolded, and the second best is underlined.
+On each task, the best method is bolded, and the second (or the third closest to the second) best is underlined.
 We use “**” to denote the method that achieves statistically better results than all the other methods (through statistical tests).
-Model Fluorescence ($\rho$uparrow) Stability ($\rho$uparrow) $\beta$-lactamase ($\rho$uparrow) Solubility (PR-AUC $\uparrow$)
+Model Fluorescence Stability $\beta$-lactamase Solubility
+$\rho$uparrow MAE $\downarrow$ $\rho$uparrow MAE $\downarrow$ $\rho$uparrow MAE $\downarrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
 # train/valid/test 21446 / 5362 / 27217 53571 / 2512 / 12851 4158 / 520 / 520 62478 / 6942 / 1999
-CNN 0.680 $\pm$ 0.001 0.715 $\pm$ 0.025** 0.721 $\pm$ 0.020** $74.61\pm 0.55$
-CNN-RNN $0.678\pm 0.001$ $0.638\pm 0.021$ $0.695\pm 0.012$ $74.10\pm 1.48$
-Transformer $0.648\pm 0.001$ $0.442\pm 0.030$ $0.348\pm 0.002$ 78.86 $\pm$ 0.46**
-GCN $0.397\pm 0.002$ $0.392\pm 0.027$ $0.417\pm 0.001$ $69.26\pm 0.73$
-GAT $0.251\pm 0.002$ $0.165\pm 0.026$ $0.196\pm 0.013$ $62.44\pm 0.01$
-NeuralFP $0.413\pm 0.011$ $0.333\pm 0.041$ $0.133\pm 0.020$ $78.74\pm 0.24$
-AttentiveFP $0.260\pm 0.003$ $0.254\pm 0.022$ $0.058\pm 0.011$ $78.39\pm 0.19$
-MPNN $0.237\pm 0.020$ $0.110\pm 0.021$ $0.068\pm 0.015$ $62.53\pm 0.31$
-PAGTN $0.188\pm 0.036$ $0.266\pm 0.016$ $0.092\pm 0.018$ $61.33\pm 0.91$
-Graphormer $0.067\pm 0.002$ OOM $0.103\pm 0.018$ OOM
+CNN 0.680 $\pm$ 0.001 0.194 $\pm$ 0.001 0.715 $\pm$ 0.025 0.312 $\pm$ 0.008 0.721 $\pm$ 0.020 0.152 $\pm$ 0.001 $74.61\pm 0.55$ 62.50 $\pm$ 2.69
+CNN-RNN $0.678\pm 0.001$ 0.268 $\pm$ 0.002 $0.635\pm 0.025$ 0.323 $\pm$ 0.001 $0.695\pm 0.012$ 0.157 $\pm$ 0.002 $75.46\pm 0.03$ 60.87 $\pm$ 3.19
+Transformer $0.648\pm 0.001$ 0.371 $\pm$ 0.005 $0.375\pm 0.052$ 0.411 $\pm$ 0.026 $0.310\pm 0.041$ 0.232 $\pm$ 0.007 78.86 $\pm$ 0.46 61.28 $\pm$ 1.42
+GCN $0.397\pm 0.002$ 0.864 $\pm$ 0.018 0.443 $\pm$ 0.017 0.445 $\pm$ 0.011 $0.450\pm 0.006$ 0.225 $\pm$ 0.001 $69.26\pm 0.73$ 56.30 $\pm$ 1.39
+GAT $0.249\pm 0.001$ 1.201 $\pm$ 0.005 $0.101\pm 0.001$ 0.740 $\pm$ 0.002 $0.196\pm 0.013$ 0.264 $\pm$ 0.002 $62.44\pm 0.01$ 47.79 $\pm$ 0.42
+NeuralFP $0.349\pm 0.049$ 1.016 $\pm$ 0.081 $0.373\pm 0.023$ 0.484 $\pm$ 0.020 $0.171\pm 0.007$ 0.263 $\pm$ 0.001 $78.74\pm 0.24$ 62.72 $\pm$ 1.91
+AttentiveFP $0.180\pm 0.006$ 1.285 $\pm$ 0.001 $0.013\pm 0.004$ 0.763 $\pm$ 0.003 $0.058\pm 0.011$ 0.262 $\pm$ 0.007 60.56 $\pm$ 0.74 39.21 $\pm$ 0.66
+MPNN $0.249\pm 0.002$ 1.275 $\pm$ 0.011 $0.118\pm 0.061$ 0.732 $\pm$ 0.006 $0.068\pm 0.015$ 0.262 $\pm$ 0.003 $62.53\pm 0.31$ 40.56 $\pm$ 5.46
+PAGTN $0.139\pm 0.014$ 1.255 $\pm$ 0.040 $0.088\pm 0.194$ 0.727 $\pm$ 0.011 $0.092\pm 0.018$ 0.255 $\pm$ 0.007 $61.33\pm 0.91$ 46.52 $\pm$ 4.13
+Graphormer $0.058\pm 0.015$ 1.201 $\pm$ 0.032 OOM OOM $0.067\pm 0.046$ 0.287 $\pm$ 0.007 OOM OOM
+ESM-1b 0.562 $\pm$ 0.003 0.548 $\pm$ 0.004 0.682 $\pm$ 0.077 0.282 $\pm$ 0.010 0.664 $\pm$ 0.002 0.171 $\pm$ 0.001 Length Exceeded Length Exceeded
+ESM-2-650M 0.548 $\pm$ 0.004 0.563 $\pm$ 0.002 0.689 $\pm$ 0.020 0.273 $\pm$ 0.014 0.694 $\pm$ 0.001 0.154 $\pm$ 0.001 71.53 $\pm$ 0.58 63.33 $\pm$ 3.68
+Prot-T5-XL 0.553 $\pm$ 0.003 0.535 $\pm$ 0.006 $\bf 0.767\pm 0.008$ 0.272 $\pm$ 0.030 0.756 $\pm$ 0.003 0.136 $\pm$ 0.003 74.95 $\pm$ 1.01 62.75 $\pm$ 0.76
+ProtBert 0.566 $\pm$ 0.001 0.583 $\pm$ 0.007 0.646 $\pm$ 0.064 0.322 $\pm$ 0.026 0.572 $\pm$ 0.001 0.207 $\pm$ 0.001 70.70 $\pm$ 0.47 62.53 $\pm$ 2.67
+DeepProt-T5 0.614 0.642 0.725 0.551 0.874** 0.084** 77.00 68.25**
+ChemLLM-7B -0.019 19.878 -0.187 46273.897 0.017 111.76 50.65 1.64
+LlaSMol-Mistral-7B NaN 2.083 0.094 1.359 NaN 0.268 50.30 6.89
 Table 3: Results of protein localization prediction.
-Model Subcellular (Acc $\uparrow$) Binary (PR-AUC $\uparrow$)
+Model Subcellular Binary
+Acc $\uparrow$ Averaged F1 $\uparrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
 # train/valid/test 8,945 / 2248 / 2768 5161 / 1727 / 1746
-CNN 50.18 $\pm$ 1.21 $90.88\pm 0.31$
-CNN-RNN 52.58 $\pm$ 0.11** 91.43 $\pm$ 0.45**
-Transformer $42.63\pm 0.68$ $78.38\pm 0.25$
-GCN $47.45\pm 0.47$ $83.43\pm 0.10$
-GAT $45.14\pm 0.10$ $82.15\pm 0.41$
-NeuralFP $45.20\pm 0.49$ $81.14\pm 0.06$
-AttentiveFP $42.38\pm 1.25$ $80.58\pm 0.30$
+CNN 50.18 $\pm$ 1.21 30.44 $\pm$ 0.28 $90.88\pm 0.31$ 87.74 $\pm$ 0.08
+CNN-RNN 52.58 $\pm$ 0.11 38.21 $\pm$ 0.57 91.43 $\pm$ 0.45 88.56 $\pm$ 1.03
+Transformer $42.63\pm 0.68$ 25.46 $\pm$ 0.01 $78.38\pm 0.25$ 72.20 $\pm$ 2.21
+GCN $47.45\pm 0.47$ 34.88 $\pm$ 0.79 $83.43\pm 0.10$ 81.80 $\pm$ 0.07
+GAT $45.14\pm 0.10$ 27.03 $\pm$ 0.47 $82.15\pm 0.41$ 81.89 $\pm$ 0.19
+NeuralFP $45.20\pm 0.49$ 27.07 $\pm$ 0.93 $81.14\pm 0.06$ 79.95 $\pm$ 0.09
+AttentiveFP $42.38\pm 1.25$ 23.50 $\pm$ 1.34 $80.58\pm 0.30$ 80.08 $\pm$ 0.39
+ESM-2-650M 79.07 $\pm$ 0.05 $6$$\pm$ 1.14 96.63 $\pm$ 0.18 91.76 $\pm$ 0.47
+Prot-T5-XL 80.67 $\pm$ 0.04 69.86 $\pm$ 0.39 97.03 $\pm$ 0.13 93.48 $\pm$ 0.07
+LlaSMol-Mistral-7B 15.65 4.99 57.52 0.00
+ChemLLM-7B 6.24 0.59 56.32 1.43
+DeepProt-T5 82.69 82.52 92.17 92.18
 Table 4: Results of Protein-Protein Interaction (PPI).
-Model PPI Affinity ($R^{2}$uparrow) Yeast (PR-AUC $\uparrow$) Human PPI (PR-AUC $\uparrow$)
+Model PPI Affinity Yeast PPI Human PPI
+$R^{2}$uparrow MAE $\downarrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
 # train/valid/test 2127 / 212 / 343 1668 / 131 / 373 6844 / 277 / 227
-CNN $0.493\pm 0.015$ $51.93\pm 0.92$ $70.37\pm 1.22$
-CNN-RNN 0.584 $\pm$ 0.026** $53.28\pm 0.85$ $70.45\pm 2.68$
-Transformer $0.425\pm 0.021$ $53.79\pm 1.07$ $59.36\pm 4.00$
-GCN $0.394\pm 0.006$ 58.98 $\pm$ 0.72 82.21 $\pm$ 1.13
-GAT $0.230\pm 0.001$ $53.72\pm 0.39$ $77.63\pm 3.13$
-NeuralFP $0.100\pm 0.054$ $57.00\pm 1.51$ $80.11\pm 1.25$
+CNN $\bf 0.646\pm 0.003$ 1.764 $\pm$ 0.051 $51.93\pm 0.92$ 25.90 $\pm$ 0.10 $70.37\pm 1.22$ 69.65 $\pm$ 1.63
+CNN-RNN 0.584 $\pm$ 0.026 1.886 $\pm$ 0.108 $53.28\pm 0.85$ 47.20 $\pm$ 3.62 $70.45\pm 2.68$ 69.87 $\pm$ 2.59
+Transformer $0.425\pm 0.021$ 2.081 $\pm$ 0.133 $53.79\pm 1.07$ 51.93 $\pm$ 0.40 $59.36\pm 4.00$ 68.64 $\pm$ 1.06
+GCN $0.366\pm 0.034$ 2.443 $\pm$ 0.036 58.98 $\pm$ 0.72 48.13 $\pm$ 3.88 82.21 $\pm$ 1.13 71.05 $\pm$ 2.90
+GAT $0.230\pm 0.001$ 2.463 $\pm$ 0.015 $53.72\pm 0.39$ 57.00 $\pm$ 3.83 $77.63\pm 3.13$ 73.92 $\pm$ 3.50
+NeuralFP $0.100\pm 0.054$ 2.555 $\pm$ 0.040 $57.00\pm 1.51$ 58.94 $\pm$ 4.74 $80.11\pm 1.25$ 67.62 $\pm$ 1.03
+ESM-2-650M 0.592 $\pm$ 0.001 1.893 $\pm$ 0.005 67.36 $\pm$ 0.80 63.99 $\pm$ 1.00 96.17 $\pm$ 0.18 87.63 $\pm$ 0.60
+Prot-T5-XL 0.573 $\pm$ 0.011 1.979 $\pm$ 0.007 69.84 $\pm$ 0.46 66.97 $\pm$ 0.01 95.36 $\pm$ 0.07 87.64 $\pm$ 0.98
+LlaSMol-Mistral-7B -0.008 20.335 53.64 51.39 49.48 49.15
+ChemLLM-7B 0.082 28.771 53.65 15.45 47.15 15.49
+DeepProt-T5 0.643 1.870 63.20 66.70 76.37 80.28
 Table 5: Results of epitope and paratope prediction (residue-level classification).
-Model IEDB (ROC-AUC $\uparrow$) PDB-Jespersen (ROC-AUC $\uparrow$) SAbDab-Liberis (ROC-AUC $\uparrow$)
+Structure-based and pretrained protein language models took large GPU or CPU memory so we disregrad them in residue-level prediction.
+The same strategy is applied to secondary structure as well.
+Model IEDB PDB-Jespersen SAbDab-Liberis
+ROC-AUC $\uparrow$ Averaged F1 $\uparrow$ ROC-AUC $\uparrow$ Averaged F1 $\uparrow$ ROC-AUC $\uparrow$ Averaged F1 $\uparrow$
 # train/valid/test 2211 / 316 / 632 313 / 45 / 89 716 / 102 / 205
-CNN $54.03\pm 0.02$ 74.46 $\pm$ 0.21** 90.85 $\pm$ 0.08
-CNN-RNN $55.47\pm 0.23$ $70.10\pm 0.97$ 96.75 $\pm$ 0.10**
-Transformer 59.79 $\pm$ 0.06** $60.10\pm 0.32$ $64.77\pm 0.11$
+CNN $54.03\pm 0.02$ 10.34 $\pm$ 0.02 74.46 $\pm$ 0.21** 52.43 $\pm$ 0.35 90.85 $\pm$ 0.08 61.51 $\pm$ 0.11
+CNN-RNN $55.47\pm 0.23$ 10.96 $\pm$ 0.26 $70.10\pm 0.97$ 44.82 $\pm$ 1.69 96.75 $\pm$ 0.10** 69.34 $\pm$ 0.10
+Transformer 59.79 $\pm$ 0.06** 16.79 $\pm$ 0.07 $60.10\pm 0.32$ 27.31 $\pm$ 0.56 $64.77\pm 0.11$ 26.73 $\pm$ 0.15
 Table 6: Results of antibody developability prediction (TAP and SAbDab-Chen) and CRISPR repair outcome prediction (CRISPR-Leenay).
-Model TAP (MAE $\downarrow$) SAbDab-Chen (MAE $\downarrow$) CRISPR-Leenay (MAE $\downarrow$)
+Model TAP SAbDab-Chen CRISPR-Leenay
+$R^{2}$uparrow MAE $\downarrow$ $R^{2}$uparrow MAE $\downarrow$ $R^{2}$uparrow MAE $\downarrow$
 # train/valid/test 169 / 24 / 48 1686 / 241 / 482 1065 / 152 / 304
-CNN $3.217\pm 0.026$ 0.219 $\pm$ 0.006 0.0745 $\pm$ 0.0005
-CNN-RNN 0.712 $\pm$ 0.069** $0.226\pm 0.001$ $0.0755\pm 0.0010$
-Transformer $3.476\pm 0.004$ $0.238\pm 0.012$ $0.1216\pm 0.0020$
-GCN $2.761\pm 0.155$ $0.326\pm 0.015$ $0.1274\pm 0.0019$
-GAT $2.675\pm 0.022$ $0.310\pm 0.010$ $0.1232\pm 0.0001$
-NeuralFP $3.436\pm 0.015$ $0.253\pm 0.011$ $0.1243\pm 0.0001$
+CNN 0.469 $\pm$ 0.066 $3.217\pm 0.026$ 0.547 $\pm$ 0.014 0.219 $\pm$ 0.006 0.781 $\pm$ 0.009 0.0745 $\pm$ 0.0005
+CNN-RNN 0.972 $\pm$ 0.008** 0.712 $\pm$ 0.069** 0.486 $\pm$ 0.012 $0.226\pm 0.001$ 0.771 $\pm$ 0.004 $0.0755\pm 0.0010$
+Transformer 0.030 $\pm$ 0.011 $3.476\pm 0.004$ 0.452 $\pm$ 0.011 $0.238\pm 0.012$ 0.200 $\pm$ 0.054 $0.1216\pm 0.0020$
+GCN 0.614 $\pm$ 0.054 $2.761\pm 0.155$ 0.434 $\pm$ 0.030 $0.326\pm 0.015$ 0.066 $\pm$ 0.010 $0.1274\pm 0.0019$
+GAT 0.777 $\pm$ 0.007 $2.675\pm 0.022$ 0.356 $\pm$ 0.015 $0.310\pm 0.010$ 0.054 $\pm$ 0.010 $0.1232\pm 0.0001$
+NeuralFP 0.205 $\pm$ 0.035 $3.436\pm 0.015$ 0.452 $\pm$ 0.027 $0.253\pm 0.011$ 0.177 $\pm$ 0.029 $0.1243\pm 0.0001$
+ESM-2-650M 0.866 $\pm$ 0.007 2.452 $\pm$ 0.027 0.600 $\pm$ 0.001 0.224 $\pm$ 0.001 0.110 $\pm$ 0.011 0.1218 $\pm$ 0.0010
+Prot-T5-XL 0.837 $\pm$ 0.010 2.417 $\pm$ 0.056 0.596 $\pm$ 0.011 0.229 $\pm$ 0.003 0.236 $\pm$ 0.004 0.1186 $\pm$ 0.0006
+LlaSMol-Mistral-7B NaN 48.104 NaN 0.241 NaN 0.7922
+ChemLLM-7B 0.099 45.957 -0.006 32.33 0.061 20.0503
+DeepProt-T5 0.758 2.922 0.536 0.186 0.753 0.078
+Table 7: Results of protein folding prediction (protein-level and residue-level classification).
+Model Fold Secondary Structure
+Acc $\uparrow$ Averaged F1 $\uparrow$ Acc $\uparrow$ Averaged F1 $\uparrow$
+# train/valid/test 12312 / 736 / 718 8678 / 2170 / 513
+CNN 8.01 $\pm$ 0.44 0.47 $\pm$ 0.11 99.44 $\pm$ 0.00 87.76 $\pm$ 0.06
+CNN-RNN 7.50 $\pm$ 0.01 1.06 $\pm$ 0.01 99.93 $\pm$ 0.00 98.80 $\pm$ 0.00
+Transformer 5.34 $\pm$ 0.24 0.20 $\pm$ 0.02 98.64 $\pm$ 0.00 49.66 $\pm$ 0.00
+GCN 8.09 $\pm$ 0.36 2.34 $\pm$ 0.15 / /
+ESM-2-650M 49.80 $\pm$ 0.36 25.55 $\pm$ 0.61 / /
+Prot-T5-XL 50.28 $\pm$ 0.52 24.24 $\pm$ 0.48 / /
+LlaSMol-Mistral-7B 0.00 0.00 / /
+ChemLLM-7B 0.64 0.04 / /
+DeepProt-T5 74.00** 73.61** / /
 3.4 Results & Analysis
-For each method, we used five different random seeds to conduct independent runs and reported the average results and their standard deviations.
-The results of protein function prediction are reported in Table 2.
+For each method, we used three different random seeds to conduct independent runs and reported the average results and their standard deviations.
+The results of protein function prediction are reported in Table 2 and Table 3.
+The results of protein-protein interaction are reported in Table 4.
+The results of epitope and paratope interaction are reported in Table 5.
+The results of antibody developability prediction are reported in Table 6.
+The results of protein structure prediction are reported in Table 7.
 Statistical Test.
@@
 We use “**” to denote the method that achieves statistically better results than all the other methods (pass statistical tests).
-Key Observations.
-We summarize the following key observations as takeaways.
-  • Sequence-based neural architectures are powerful compared with graph-based neural architectures.
-Sequence-based neural architectures, such as CNN, RNN, and transformer, obtain superior performance in most protein sequence learning tasks.
-Specifically, in 12 out of all the 15 tasks across various protein sequence learning tasks, the sequence-based model (CNN, RNN, and transformer) takes the top-2 position.
-  • Among all the 12 GNN-solvable tasks (except residue-level classification), graph neural networks (GNN) obtain the best accuracy among all the methods only in two protein-protein interaction (PPI) tasks, including Yeast and Human PPI.
-  • Among all the graph neural networks (GNNs) across the whole 12 GNN-solvable tasks (except residue-level classification), the earliest variant, GCN (Kipf and Welling, 2016), achieves the best performance in 9 tasks.
+Figure 4: We recorded training time and GPU memory assumptions for each task and each model.
+Specifically, we extract three representative methods: CNN, GCN and Prot-T5 from sequence-based, structure-based and pre-trained protein language models.
+We observed that Prot-T5 and GCN took up more GPU memory than CNN.
+Prot-T5, where upstream embeddings are fixed, is more efficient in training downstream tasks.
+Training a GCN model took more time than training a CNN or a Prot-T5 model.
+Key Observations. We summarize the following key observations as takeaways.
+  • pre-trained protein language models and our DeepProt-T5 are powerful compared with sequence-based and structure-based neural architectures.
+Sequence-based neural architectures, such as CNN, RNN, and transformer, obtain also superior performance in most protein sequence learning tasks.
+Specifically, in 12 out of all the 17 tasks across various protein sequence learning tasks, both sequenced-based models (CNN, RNN, Transformer) and the pre-trained protein language models (Prot-T5-XL, ESM-2-650M and DeepProt-T5) takes the top-2 position.
+  • Among all the 13 GNN-solvable tasks (except residue-level classification), graph neural networks (GNN) obtain the inferior performance compared with sequenced-based and protein language models.
+The potential reason would be that SMILES or original string didnt provide the 3d information (coordinates) about a protein, the graph topology given by edge featurizer is ill-defined in tbe deep graph library.
+  • Among all the graph neural networks (GNNs) across the whole 12 GNN-solvable tasks (except residue-level classification), the earliest variant, GCN (Kipf & Welling, 2016), achieves the best performance in 9 tasks.
   • Stability.
-From the learning curve (Figure 7-10), we find that GNN’s training curve is not stable.
+From the learning curve (Figure 5), we find that GNN’s training curve is not stable.
 In contrast, the sequence-based models, including CNN, RNN, and transformer, converge more stably from the learning curve.
-This can be observed from Figure 7-9.
+This can be observed from Figure 5.
 On the contrary, training is more stable, fast and accurate for GAT when it comes to TAP dataset.
   • Computational complexity.
-The runtime and memory costs are reported in Figure 35 and 6.
+The runtime and memory costs are reported in Figure 4.
 We find that GNN-based models are typically computationally inefficient.
@@
 While another model, such as CNN, RNN, and transformer, uses amino acid sequences as the input feature.
-Figure 3: Empirical computational complexity on CRISPR and SAbDab Chen measured by runtime and memory.
 4 Conclusion
@@
 As the field of proteomics continues to evolve, DeepProtein stands to contribute substantially to our understanding of protein functions, localization, and interactions, ultimately driving forward discoveries that can impact biotechnology and medicine.
-Appendix A Hyperparameter Settings
+Figure 5: Training Loss of selected datasets: PPI_Affinity, SAbDab_Chen, TAP, and SubCellular.
+Appendix A Evaluation Metrics
+In this section, we describe the basic evaluation metrics for both classification and regression tasks. In the part optimization flow it would be further detailed on the end to end training flow.
+Classification metrics.
+Most classification tasks are binary classification, except subcellular prediction in protein localization prediction, which is a 10-category classification problem, where we use accuracy (acc) (the fraction of correctly predicted/classified samples) as the evaluation metric.
+In binary classification, there are four kinds of test data points based on their ground truth and the model’s prediction,
+  1. 1.
+    positive sample and is correctly predicted as positive, known as True Positive (TP);
+  2. 2.
+    negative samples and is wrongly predicted as positive samples, known as False Positive (FP);
+  3. 3.
+    negative samples and is correctly predicted as negative samples, known as True Negative (TN);
+  4. 4.
+    positive samples and is wrongly predicted as negative samples, known as False Negative (FN).
+  • Precision. The precision is the performance of a classifier on the samples that are predicted as positive. It is formally defined as $\text{precision}=\textit{TP}/(\textit{TP}+\textit{FP}).$
+  • Recall. The recall score measures the performance of the classifier to find all the positive samples. It is formally defined as $\text{recall}=\textit{TP}/(\textit{TP}+\textit{FN})$.
+  • PR-AUC (Precision-Recall Area Under Curve).
+The area under the Precision-Recall curve summarizes the trade-off between the true positive rate and the positive predictive value for a predictive model using different probability thresholds.
+  • ROC-AUC Area Under the Receiver Operating Characteristic Curve summarizes the trade-off between the true positive rate and the false positive rate for a predictive model using different probability thresholds.
+ROC-AUC is also known as the Area Under the Receiver Operating Characteristic curve (AUROC) in some literature.
+For all these metrics, the numerical values range from 0 to 1, a higher value represents better performance.
+Regression metrics. In the regression task, both ground truth and prediction are continuous values.
+  • Mean Squared Error (MSE) measures the average of the squares of the difference between the forecasted value and the actual value.
+It is defined as $\text{MSE}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-{\hat{y}}_{i})^{2},$ where $N$ is the size of the test set; $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively.
+MSE value ranges from 0 to positive infinity.
+A lower MSE value indicates better performance.
+  • Mean Absolute Error (MAE) measures the absolute value of the difference between the predicted value and the actual value.
+It is defined as $\text{MAE}=\frac{1}{N}\sum^{N}_{i=1}|y_{i}-{\hat{y}}_{i}|,$ where $N$ is the size of the test set; $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively.
+MAE value ranges from 0 to positive infinity.
+It emphasizes the ranking order of the prediction instead of the absolute value.
+A lower MAE value indicates better performance.
+  • Spearman rank correlation ($\bf{\rho}$), also known as Spearman’s $\rho$, is a nonparametric statistical test that measures the association between two ranked variables.
+A higher $\rho$ value indicates better performance.
+  • R-squared ($R^{2}$) score is defined as the proportion of the variation in the dependent variable that is predictable from the independent variable(s).
+It is also known as the coefficient of determination in statistics.
+Higher $R^{2}$ scores indicate better performance.
+Appendix B Optimization Flow
+Dataset Selection and Processing Flow
+As mentioned in the introduction part and Table 1, previous benchmarks either lack 1) the state-of-the-art deep learning methods 2) the diverse real-world data 3) easy-to-use files for researchers outside the computer science domain to use.
+Hence we collected the data from two main databases which contain approximately 20+ protein tasks which are enough for downstream testing.
+From them we deleted the tasks that were related to the drug, especially the task drug-target interaction since the DeepPurpose library supported such functionality.
+For the datasets in the PEER benchmark, DeepProtein just inherited the functions which transformed the data into standard torch datasets, and for the TDC data they were transformed to the standard torch dataset similarly.
+When loading the dataset, it will load a pair of (protein sequence, aim) or a triple of (protein sequence 1, protein sequence 2, aim), depending on the task type.
+Featurization Flow
+Since we are talking about training here instead of inference, we ignore the featurization flow of large language models.
+We mainly consider three types of methods here, which are sequence-based, structure-based and pretraind protein language models.
+Sequence-based models take the tokenized SMILES string as the input X,
+CNN: $\quad X^{(l)}=X^{(l-1)}*W^{(l)}+b^{(l)}$ (1)
+RNN: $\quad h_{t}=\sigma(W_{h}X_{t}+U_{h}h_{t-1}+b_{h})$
+Attention: $\quad\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
+In CNN, $W^{(l)}$ is the weight matrix at layer l which is convoluted by the last hidden layer input $X^{(l-1)}$ and $b^{(l)}$ is the bias.
+The hidden state is decided by $X^{(l)}$.
+In RNN, we take each token in $X$ as the input at each time step t, where $W_{h}$ is the weight matrix for the current input token (amino acid) $X_{t}$ and $U_{h}$ is the weight matrix for the last hidden state and $b_{h}$ is the bias.
+We note that the protein sequence would be long in real-world data, so we truncate them to the maximum length 300, which also avoids memory exploding in RNN.
+In transformer, for each attention block, we could compute Q, K, V by $W_{Q}X$, $W_{K}X$, and $W_{V}X$, then attention is computed by equation (1).
+Noted that we could aggregated heads of attention to perform multi-head attention.
+Structure-based models take the graph $\mathcal{G}$ as the input, with node features $\mathbf{H}^{(0)}$ and adjacency matrix $\mathbf{A}$.
+Edge features could be added well if it’s well prepared by the dataset.
+Especailly for the 2D protein structure, we could only obtain node features by using the features from CNN for instance.
+GCN, GAT and Graph Transformer’s forms are given by:
+GCN: $\quad\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$ (2)
+GAT: $\quad\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{j\in\mathcal{N}(i)\cup\{i\}}\alpha_{ij}^{(l)}\mathbf{W}^{(l)}\mathbf{h}_{j}^{(l)}\right),$
+$\quad\alpha_{ij}^{(l)}=\frac{\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}\|\mathbf{W}^{(l)}\mathbf{h}_{j}^{(l)}\right]\right)\right)}{\sum_{k\in\mathcal{N}(i)\cup\{i\}}\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}\|\mathbf{W}^{(l)}\mathbf{h}_{k}^{(l)}\right]\right)\right)}$
+Graph Transformer: $\quad\mathbf{H}^{(l+1)}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{A}\right)\mathbf{V},$
+$\quad\mathbf{Q}=\mathbf{H}^{(l)}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{H}^{(l)}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{H}^{(l)}\mathbf{W}_{V}$
+MPNN: $\quad\mathbf{m}_{v}^{(l)}=\sum_{u\in\mathcal{N}(v)}M\left(\mathbf{h}_{v}^{(l)},\mathbf{h}_{u}^{(l)},\mathbf{e}_{vu}\right),$
+$\quad\mathbf{h}_{v}^{(l+1)}=U\left(\mathbf{h}_{v}^{(l)},\mathbf{m}_{v}^{(l)}\right)$
+In GCN, $\mathbf{H}^{(l)}$ is the node representation at layer l, $\tilde{\mathbf{A}}$ is the adjacency matrix with self loop added.
+$\tilde{\mathbf{D}}$ is the degree matrix corresponding to $\tilde{\mathbf{A}}$.
+$\mathbf{W}^{(l)}$ is the weight matrix at layer l.
+in GAT, $\alpha_{ij}^{(l)}$ works as a trainable attention parameter to present the attention betwen node i and node j at layer l.
+In a general graph transformer, adjacency matrix is added to the attention term which is different from the vanilla self-attention block.
+Therefore, the complexity is still $O(n^{2})$ if there’re $n^{2}$ nodes in the graph.
+In the message passing neural network (MPNN), additional edge information $e_{vu}$ is considered for node v and node u.s
+For the pre-trained protein languade models (PLM), the general form could be written as:
+X=PLM($\textbf{X})$ = PLM ( X ) (3)
+where we regard PLM as a white box model.
+We could get the embedding for the whole protein sequence instead of encoding each amino acid one by one which is more efficient than sequence-based or structure-based encoding.
+Training Flow
+After we obtained features from the featurizer module, we train the downstream tasks with a linear layer with the weight W and bias $b$.
+We consider five machine learning task types (not referring to protein learning tasks), which are single protein regression, single protein classification, protein pair regression, protein pair classification, and token (residue) level single protein classification.
+We introduce them one by one.
+Single protein regression task is that given a single protein’s representation X, after applying the linear layer, we got a floating-point number $\hat{y}$, so mean squared error loss between the true value $y_{i}$ and predicted value $\hat{y}_{i}$ is applied during training.
+For single protein classification task, we apply the softmax function after the linear layer to decide its class.
+Either a binary cross-entropy loss (BCELoss) or a general cross-entropy loss (CELoss) would be back-propagated during the training:
+[…]
+$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$
+[…]
+$\mathcal{L}_{\text{CE}}=-\sum_{i}y_{i}\log(\hat{y}_{i})$
+Protein pair regression is that, the input is a pair of protein ($X_{i},X_{j}$), the aim is to predict its affinity or some other related interaction metrics, labeled $y_{ij}$ here.
+The representation of two proteins is concatenated before it is applied to a linear layer.
+The predicted value is $\hat{y}_{ij}$.
+MSE loss between $\hat{y}_{ij}$ and $y_{ij}$ would be back-propagated.
+For protein pair classification task, we apply a sigmoid function as all labels are either 0 or 1 in our benchmark.
+BCELoss is being computed.
+[…]
+$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i,j}(\hat{y}_{ij}-y_{ij})^{2}$
+[…]
+$\mathcal{L}_{\text{BCE}}=-\sum_{i,j}y_{ij}\log(\hat{y}_{ij})+(1-y_{ij})\log(1-\hat{y}_{ij})$
+For residue-level single protein classification, we predict the class for each token (amino acid) for each protein sequence, from the token $X_{1}$ to the the token $X_{T}$ if the length is equal to T.
+A softmax is applied here after applying a linear layer and CE-loss is calculated.
+Note that it is computation inefficient to perform residue level prediction when applying graph neural networks and protein language models and could easily reach memory bound so in our benchmark we only tested those datasets with CNN, CNN-RNN and Transformer architectures.
+$\hat{Y}_{t,c}$ is the probability that $X_{t}$ is being assigned to class c.
+[…]
+$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{c}Y_{t,c}\log(\hat{Y}_{t,c})$
+Appendix C Tables of Time and Memory Usage
+Table 8: Memory and Time Usage of Different Models
+Dataset Model Time (hour) GPU Memory (GB)
+Fluorescence CNN 0.01000 0.78000
+Fluorescence GCN 0.13200 1.00000
+Fluorescence Prot-T5 0.00160 5.20000
+Stability CNN 0.01600 0.84480
+Stability GCN 0.06100 0.73656
+Stability Prot-T5 0.00472 10.19000
+Beta CNN 0.00150 0.84700
+Beta GCN 0.14000 1.07000
+Beta Prot-T5 0.00060 5.20000
+Solubility CNN 0.01300 0.84700
+Solubility GCN 0.48000 2.70200
+Solubility Prot-T5 0.04000 7.33440
+SubCellular CNN 0.00600 2.94500
+SubCellular GCN 0.07000 8.72600
+SubCellular Prot-T5 0.00060 5.71200
+SubCellular-Bin CNN 0.00110 1.82600
+SubCellular-Bin GCN 0.05200 10.75200
+SubCellular-Bin Prot-T5 0.00030 5.62300
+PPI_Affinity CNN 0.00400 1.37000
+PPI_Affinity GCN 0.03000 1.64600
+PPI_Affinity Prot-T5 0.00050 5.16200
+Yeast-PPI CNN 0.01000 2.69000
+Yeast-PPI GCN 0.32000 12.62000
+Yeast-PPI Prot-T5 0.00020 9.04000
+Human-PPI CNN 0.14000 0.96000
+Human-PPI GCN 1.10000 2.13000
+Human-PPI Prot-T5 0.00320 9.38000
+TAP CNN 0.01000 1.95000
+TAP GCN 0.00300 0.69100
+TAP Prot-T5 0.00030 5.16000
+SAbDab-Chen CNN 0.00300 0.65000
+SAbDab-Chen GCN 0.01000 1.05000
+SAbDab-Chen Prot-T5 0.00060 5.16000
+CRISPR CNN 0.00020 0.62900
+CRISPR GCN 0.00100 0.53800
+CRISPR Prot-T5 0.00050 5.16000
+Fold CNN 0.01000 0.92200
+Fold GCN 0.06000 1.80700
+Fold Prot-T5 0.00020 5.22200
+Appendix D Prompt Template
+Template for ChemLLM-7B:
+System Instruction: You are an AI assistant specializing in protein property prediction. Follow the given instruction format. User Prompt Format: <|im_start|>user
+What is the {protein_property} of the given protein sequence {protein_sequence}? {instruction}
+<|im_end|>
+<|im_start|>assistant
+Template for LlaSMol-Mistral-7B:
+[…]
+Instruction and property (task) for each dataset.
+Datasets Instruction Property
+Fluorescence You should return a floating-point number. Fluorescence intensity
+Beta You should return a floating-point number. Increased activity
+Stability You should return a floating-point number. Protein stability
+Solubility You should return an integer (0 or 1) where 0 is not soluble and 1 is soluble. Protein solubility
+Subcellular You should choose an integer within the range [0, 9] to indicate the protein’s location. Location
+Subcellular_Binary You should return an integer (0 or 1) where 0 is membrane-bound and 1 is soluble. Location
+Tap You should return a floating-point number. Developability
+SAbDab_Chen You should return a floating-point number. Developability
+CRISPR You should return a floating-point number. Repair outcome
+PPI-Affinity You should return a floating-point number. Activity of protein-protein interaction
+Yeast-PPI You should return an integer (0 or 1) where 0 is weak and 1 is strong. Activity of protein-protein interaction
+Human-PPI You should return an integer (0 or 1) where 0 is weak and 1 is strong. Activity of protein-protein interaction
+Fold You should return an integer within the range [0, 1194]. Global structural topology of a protein on the fold level
+Secondary You should return an integer within the range [0, 2]. Local structures of protein residues in their natural state
+Appendix E Hyperparameter Settings
 In table 10, we have listed a common settings of hyperparameter used in this library.
@@
 For Subcellular and its binary version, a training epoch of 60 is enough for convergence.
-For small-scale protein datasets such as IEDB [Yi et al., 2018], PDB-Jespersen, and SAbDab-Liberis, a larger learning rate of 0.001 also leads to convergence and the same performance when using CNN, CNN-RNN, and Transformer.
+For small-scale protein datasets such as IEDB (Yi et al., 2018), PDB-Jespersen, and SAbDab-Liberis, a larger learning rate of 0.001 also leads to convergence and the same performance when using CNN, CNN-RNN, and Transformer.
 For TAP, SAbDab-Chen and CRISPR-Leenay, larger learning rate of 0.0001 is suggested when training graph neural networks.
@@
 Graphormer $10^{-5}$ 0.1 ReLU 8 1 64 MaxPooling 32 100 LayerNorm
-Appendix B More results of performance on protein sequence learning
-In this section, we show three more model performance on Fluorescence, Beta-lactamase and Human PPI dataset.
-Figure 4: Performance of selected models including CNN, Transformer, GCN and NeuralFP on Fluorescence, Beta-lactamase and Human PPI.
-We observe that CNN performs significantly better than GCN on protein function prediction tasks.
-However, GCN dominates on large-scale PPI datasets such as the Human PPI dataset.
-Appendix C More results of empirical memory and runtime complexity
-In this section, we show four more empirical complexity results on Fluorescence, Beta-lactamase, SubCellular Binary and Yeast PPI.
-In Figure 5, the same pattern is found as Figure 3: Transformer took up a large amount of memory to train while the training speed is reasonably fast.
-Training GNN is slow since the DGL framework computes the message passing on the CPU.
-CNNs are both fast and memory efficient to train.
-A different case is found in Figure 6 where for SubCellular Binary and Yeast PPI, GCN is the slowest to train with occupying a large amount of GPU memory space.
-Figure 5: Empirical computational complexity on CRISPR and SAbDab Liberis measured by runtime and memory.
-Figure 6: Empirical computational complexity on SubCellular Binary and Yeast PPI measured by runtime and memory.
-Appendix D Training Curves of DeepProtein
-In this section, we plot the metrics (Accuracy or R2) and training loss for SubCellular, PPI Affinity, SAbDab-Chen and TAP.
-Figure 7: SubCellular
-Figure 8: PPI Affinity
-Figure 9: SAbDab_Chen
-Figure 10: TAP
-Appendix E Few Lines of Codes
-Protein Function Prediction
-We show an example of CNN on Beta_lactamase with DeepProtein.
-More case studies can be found in the GitHub repository of DeepProtein.
-We annotate the code to make researchers better understand the function of code base, which is simpler and clearer than most of the protein learning libraries available.
-$>$ import os, sys, argparse, torch, wandb
-$>$ from DeepProtein.dataset import *
-$>$ import DeepProtein.utils as utils
-$>$ import DeepProtein.ProteinPred as models
-$>$ path = os.getcwd()
-$>$ train_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’train’)
-$>$ valid_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’valid’)
-$>$ test_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’test’)
-$>$ train_protein_processed,train_target,train_protein_idx = collate_fn(train_beta)
-$>$ valid_protein_processed,valid_target,valid_protein_idx = collate_fn(valid_beta)
-$>$ test_protein_processed,test_target,test_protein_idx = collate_fn(test_beta)
-$>$ target_encoding = ’CNN’
-$>$ train, _, _ = utils.data_process(X_target=train_protein_processed, y=train_target, target_encoding=target_encoding, split_method=’random’, frac=[0.99998, 1e-5, 1e-5])
-$>$ _, val, _ = utils.data_process(X_target=valid_protein_processed, y=valid_target, target_encoding=target_encoding, split_method=’random’, frac=[1e-5, 0.99998, 1e-5])
-$>$ _, _, test = utils.data_process(X_target=test_protein_processed, y=test_target, target_encoding=target_encoding,split_method=’random’, frac=[1e-5, 1e-5, 0.99998])
-$>$ config = generate_config(target_encoding=target_encoding, \
-cls_hidden_dims=[1024, 1024], \
-train_epoch=20, \
-LR=0.0001, \
-batch_size=32, \
-$>$ config[’multi’] = False
-$>$ torch.manual_seed(args.seed)
-$>$ model = models.model_initialize(**config)
-$>$ model.train(train, val, test, compute_pos_enc = False)
-If we use argparse in Python, this can be easily done in one line, where
-python beta.py --target_encoding CNN
---seed 7
---wandb_proj DeepPurposePP
---lr 0.0001
---epochs 100
-Note that if you want to observe the result offline without wandb, then set
-wandb.init(mode=’offline’, name = "your_project")
-inside the training file.
-Appendix F Tutorial
-Demos can be found under DEMOS folder in DeepProtein.
+Appendix F Competing interests
+No competing interest is declared.
+Appendix G Author contributions statement
+J.X. and T.F. conceived the experiment(s), J.X. conducted the experiment(s), and J.X., Y.Z., and T.F. analyzed the results. J.X., Y.Z., and T.F. wrote and reviewed the manuscript.
+Appendix H Acknowledgments
+The authors thank the anonymous reviewers for their valuable suggestions. Y.Z. is partly supported by National Science Foundation (NSF) Award No. 2346158.
+Appendix I Appendix
+You may include other additional sections here.